### v1

# AdaDetectGPT: Adaptive Detection of LLM-Generated Text with Statistical Guarantees

Abstract
We study the problem of determining whether a piece of text has been authored by a human or by a large language model (LLM). Existing state of the art logits-based detectors make use of statistics derived from the log-probability of the observed text evaluated using the distribution function of a given source LLM. However, relying solely on log probabilities can be sub-optimal. In response, we introduce AdaDetectGPT – a novel classifier that adaptively learns a witness function from training data to enhance the performance of logits-based detectors. We provide statistical guarantees on its true positive rate, false positive rate, true negative rate and false negative rate. Extensive numerical studies show AdaDetectGPT nearly uniformly improves the state-of-the-art method in various combination of datasets and LLMs, and the improvement can reach up to 58%. A python implementation of our method is available at https://github.com/Mamba413/AdaDetectGPT.
1 Introduction
Large language models (LLMs) such as ChatGPT (OpenAI, 2022), PaLM (Chowdhery et al., 2023), Llama (Grattafiori et al., 2024) and DeepSeek (Bi et al., 2024) have revolutionized the field of generative artificial intelligence by enabling large-scale content generation across various fields including journalism, education, and creative writing (Demszky et al., 2023; Milano et al., 2023; Doshi & Hauser, 2024). However, their ability to produce highly human-like text poses serious risks, such as the spread of misinformation, academic dishonesty, and the erosion of trust in written communication (Ahmed et al., 2021; Lee et al., 2023; Christian, 2023). Consequently, accurately distinguishing between human- and LLM-generated text has emerged as a critical area of research.
There is a growing literature on the detection of machine-generated text; refer to Section 1.1 for a review. One popular line of research focuses on statistics-based detectors, typically rely on log-probability outputs (i.e., logits) from a source LLM to construct the statistics for classification (see e.g., Gehrmann et al., 2019; Mitchell et al., 2023). These works are motivated by the empirical observation that LLM-generated text tends to exhibit higher log-probabilities or larger differences between the logits of original and perturbed tokens. However, as we demonstrated in Section 3, relying solely on the logits can be sub-optimal for detecting LLM-generated text.
Our contribution. In this paper, we propose AdaDetectGPT (see Figure 1 for a visualization), an adaptive LLM detector that leverages external training data to enhance the effectiveness of existing logits-based detectors. Our approach derives a lower bound on the true negative rate (TNR) of logits-based detectors and adaptively learns a witness function by optimizing this bound, resulting in a more powerful detection statistic. The optimization is straightfoward and requires only solving a system of linear equations. Based on this statistic, we further introduce an approach to select the classification threshold that controls AdaDetectGPT’s false negative rate (FNR).
Empirically, we conduct extensive evaluations across multiple datasets and a variety of target language models to demonstrate that AdaDetectGPT consistently outperforms existing logits-based detectors. In white-box settings – where the target LLM to be detected is the same as the source LLM used to compute the logits – AdaDetectGPT achieves improvements over the best alternative in area under the curve (AUC) ranging from 12.5% to 37% over the best existing method. In black-box settings, where the source and target LLMs differ, it can offer gains of up to 58%.
Theoretically, we provide statistical performance guarantees for AdaDetectGPT, deriving finite-sample error bounds for its TNR, FNR, true positive rate (TPR) and false positive rate (FPR). Existing literature on logits-based detectors generally lacks systematic statistical analysis. Our work aims to fill in this gap and contribute toward a deeper understanding of these methods in this emerging field, by offering a comprehensive analysis based on the aforementioned standard classification metrics.
Figure 1: Workflow of AdaDetectGPT. Built upon Fast-DetectGPT (Bao et al., 2024), our method adaptively learn a witness function $\widehat{w}$ from training data by maximizing a lower bound on the TNR, while using normal approximation for FNR control.
1.1 Related works
Existing methods for detecting machine-generated text generally fall into three categories: machine learning (ML)-based, statistics-based, and watermarking-based; see Wu et al. (2025) for a recent comprehensive survey. We review the first two categories below and defer the discussion of watermarking-based approaches to Appendix A, since our method, unlike watermarking-based approaches, does not depend on knowledge of the specific hash function or random number generator used during token generation, which are often model-specific and not publicly accessible.
ML-based detection. ML-based methods train classification models on external human- and machine-authored text for detection. Many methods can be further categorized into two types. The first type extracts certain features from text and apply classical ML models to train classifiers based on these features. Various features have been proposed in the literature, ranging from classical term frequency-inverse document frequency (TF-IDF), unigram, and bigram features (Solaiman et al., 2019), to more complex features engineered specifically for this task, such as the cross-entropy loss computed between the source text and a surrogate LLM (Guo et al., 2024a) and the rewriting-based measure that quantifies the difference between original texts and their LLM-rewritten versions (Mao et al., 2024).
The second type of methods fine-tune LLMs directly for classification. This approach is intuitive, as LLMs are inherently designed for processing text data; we only need to modify the model’s output to predict a binary label rather than token probabilities. Various LLMs have been employed for fine-tuning, including RoBERTa (Solaiman et al., 2019; Guo et al., 2023), BERT (Ippolito et al., 2020) and DistilBERT (Mitrović et al., 2023).
In addition to these two types of methods, Abburi et al. (2023) propose a hybrid approach that uses the outputs of fine-tuned LLMs as input features for classical ML-based classification. Further efforts have focused on handling adversarial attacks (Crothers et al., 2022; Krishna et al., 2023; Koike et al., 2024; Sadasivan et al., 2025), short texts (Tian et al., 2024), out-of-distribution texts Guo et al. (2024b), biases against non-native English writers (Liang et al., 2023), as well as the downstream applications of these methods in domains such as education, social media and medicine (Herbold et al., 2023; Kumarage et al., 2023; Liao et al., 2023).
Statistics-based detection. Statistics-based methods leverage differences in token-level metrics such as log-probabilities to distinguish between human- and machine-authored text. Unlike ML-based approaches, many of these methods do not rely on external training data; instead, they directly use predefined statistical measures as classifiers. In particular, a seminal work by Gehrmann et al. (2019) propose several such measures, including the average log-probability and the distribution of absolute ranks of probabilities of tokens across a text. These measures exhibit substantial differences between human- and machine-authored text, and they have been widely employed and extended in the literature (see e.g., Mitchell et al., 2023; Su et al., 2023; Bao et al., 2024; Hans et al., 2024).
Other statistical measures employed are calculated based on the N-gram distributions (Solaiman et al., 2019; Yang et al., 2024b), the intrinsic dimensionality of text (Tulchinskii et al., 2023), the reward model used in LLMs (Lee et al., 2024) and the maximum mean discrepancy (Zhang et al., 2024; Song et al., 2025). Recent works have extended these methods to more challenging scenarios, such as to handle adversarial attacks (Hu et al., 2023) and black-box settings (Yu et al., 2024; Zeng et al., 2024). Theoretically, Chakraborty et al. (2024) establish a sample complexity bound for detecting machine-generated text.
To conclude this section, we remark that our proposal lies at the intersection between statistics- and ML-based methods. Similar to many statistics-based approaches, our classifier is constructed based on the log-probabilities. However, we adaptively learn a witness function via ML to improve its effectiveness. In this way, our method leverages the strengths of both approaches, leading to superior detection performance.
2 Preliminaries
We first define the white-box and black-box settings as well as our objective. We next review two baseline methods, DetectGPT (Mitchell et al., 2023) and Fast-DetectGPT (Bao et al., 2024), as they are closely related to our proposal. Finally, we introduce the martingale central limit theorem (see e.g., Hall & Heyde, 2014), which serves as the theoretical basis for our threshold selection.
Task and settings. We study the problem of determining whether a given passage $\bm{X}$, represented as a sequence of tokens $(X_{1},X_{2},\ldots,X_{L})$, was authored by a human or generated by a target LLM. Specifically, let $p$ and $q$ denote the distributions over human-written and LLM-generated tokens, respectively. Each distribution can be represented as a product of conditional probability functions, $\prod_{t}p_{t}$ for humans and $\prod_{t}q_{t}$ for the target LLM, where each $p_{t}(x_{t}|x_{<t})$ (and similarly $q_{t}$) denotes the conditional probability mass function of the next token $x_{t}$ given the preceding tokens, where $x_{<t}\coloneqq(x_{1},x_{2},\ldots,x_{t-1})$ when $t>1$ and $x_{<t}\coloneqq\emptyset$ otherwise. Our goal is to develop a classifier to discriminate between $\bm{X}\sim p$ (human) and $\bm{X}\sim q$ (LLM).
We assume access to a source LLM’s probability distribution function $q^{\prime}=\prod_{t}q_{t}^{\prime}$. When $q^{\prime}=q$, it corresponds to the white-box setting where the source model we have is the same as the target model we wish to detect. This is the primary setting considered in this paper. For closed-source LLMs such as GPT-3.5 and GPT-4, their probability functions are not publicly available. In such cases, we utilize an open-source model with distribution $q^{\prime}$ as an approximation of $q$, resulting in the black-box setting, which our method is also extended to handle.
We also assume access to a corpus of $n$ human-authored passages $\mathcal{H}=\{\bm{X}^{(i)}\}_{i=1}^{n}\sim p$. This assumption is reasonable, as large corpora of human-written text are readily available online (e.g., Wikipedia). Without loss of generality, we assume that all passages have the same number of tokens $L$, achieved by zero-padding shorter sequences to match the maximum token length. Throughout this paper, we use boldface letters (e.g., $\bm{X}$) to denote passages, and non-boldface letters (e.g., $X$) to denote tokens.
Baseline methods. Both DetectGPT and Fast-DetectGPT are statistics-based and rely on the log-probability of a passage, $\log q^{\prime}(\bm{X})$, as the basis for classification. Specifically, DetectGPT considers the following statistic:
$\frac{\log q^{\prime}(\bm{X})-\mathbb{E}_{\widetilde{\bm{X}}\sim p^{\prime}(\bullet|\bm{X})}[\log q^{\prime}(\widetilde{\bm{X}})]}{\sqrt{\textrm{Var}_{\widetilde{\bm{X}}\sim p^{\prime}(\bullet|\bm{X})}(\log q^{\prime}(\widetilde{\bm{X}}))}}$, (1)
where both the expectation in the numerator and the variance in the denominator are evaluated under a perturbation function $p^{\prime}$, which produces $\widetilde{\bm{X}}$ that is a slightly modified version of $\bm{X}$ with similar meaning. The rationale behind this statistic is that, empirically, machine-generated text tends to yield higher values than human-written text when evaluated using (1) (Mitchell et al., 2023, Figure 2). As a result, a passage is classified as machine-generated if this statistic is larger than a certain threshold.
A potential limitation of DetectGPT is that sampling from the perturbation distribution $p^{\prime}$ requires multiple calls to the source LLM to generate rewritten versions of the input passage, making the calculation of (1) computationally expensive. Fast-DetectGPT addresses this issue by proposing a modified version of (1), given by
$\frac{\sum_{t}\log q_{t}^{\prime}(X_{t}|X_{<t})-\sum_{t}\mathbb{E}_{\widetilde{X}_{t}\sim s_{t}(\bullet|X_{<t})}\log q_{t}^{\prime}(\widetilde{X}_{t}|X_{<t})}{\sqrt{\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim s_{t}(\bullet|X_{<t})}(\log q_{t}^{\prime}(\widetilde{X}_{t}|X_{<t}))}}.$ (2)
Specifically, notice that $\log q^{\prime}(\bm{X})$ can be decomposed as a token-wise sum $\sum_{t}\log q_{s}(X_{t}|X_{<t})$. Thus, the first term in the numerator of (1) is the same as that in (2). However, Fast-DetectGPT replaces the centering term in (1) with $\sum_{t}\mathbb{E}_{\widetilde{X}_{t}\sim s_{t}(\bullet|X_{<t})}\log q_{t}^{\prime}(\widetilde{X}_{t}|X_{<t})$. Here, $s=\prod_{t}s_{t}$ denotes a sampling distribution function which may equal $q$ or be derived from another LLM. By replacing the perturbation function with $s$, the centering term can be efficiently computed directly from the LLM’s conditional probabilities. Additionally, due to the conditioning on $X_{<t}$ in the centering term, the variance term is equal to the sum of the conditional variances of $\log q_{t}^{\prime}(\widetilde{X}_{t}|X_{<t})$. Finally, it classifies a passage as machine-generated if this modified statistic is larger than a certain threshold.
Martingale central limit theorem. The martingale central limit theorem (MCLT) is a fundamental result in probability theory that enables rigorous statistical inference for time-dependent data. It is well-suited for analyzing text data where tokens are generated sequentially given their predecessors. Consider a time series $\{Z_{t}\}_{t}$ where each $Z_{t}$ represents a real-valued random variable. Suppose there exists a sequence of monotonically increasing sets of random variables $\mathcal{F}_{1}\subseteq\mathcal{F}_{2}\subseteq\cdots$ so that $Z_{t}\in\mathcal{F}_{t}$ for any $t$. Under certain regularity conditions, MCLT states that the normalized partial sum
$\frac{\sum_{t=1}^{L}[Z_{t}-\mathbb{E}(Z_{t}|\mathcal{F}_{t-1})]}{\sqrt{\sum_{t=1}^{L}\textrm{Var}(Z_{t}|\mathcal{F}_{t-1})}}$ (3)
converges in distribution to a standard normal random variable as $L$ approaches infinity (Brown, 1971). Notice that the statistic in (3) shares similar structures with that employed in Fast-DetectGPT (see (2)). In the next section, we will leverage this connection for FNR control through normal approximation.
3 AdaDetectGPT
We present AdaDetectGPT in this section. We begin by discussing the white-box setting in Parts (a)(c): Part (a) introduces the proposed statistical measure; Part (b) discusses our choice of the classification threshold; Part (c) derives a lower bound on the TNR to learn the witness function. Next, in Part (d), we extend our proposal to the black-box setting. Finally, in Part (e), we establish the statistical properties of AdaDetectGPT.
(a) Statistical measure. Notice that $q^{\prime}=q$ under the white-box setting. Given a passage $\bm{X}$, the proposed classifier is built upon the following statistic:
$T_{w}(\bm{X})\coloneqq\frac{\sum_{t}[w(\log q_{t}(X_{t}|X_{<t}))-\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}(\bullet|X_{<t})}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))]}{\sqrt{\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}(\bullet|X_{<t})}(w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}},$ (4)
where $w:\mathbb{R}\to\mathbb{R}$ denotes a one-dimensional witness function defined over the space of logits.
By definition, $T_{w}(\bm{X})$ is very similar to Fast-DetectGPT’s statistic in (2), with two modifications: (i) First, rather than using the raw log conditional probability $q_{t}$, we apply a witness function $w$ to these $\log q_{t}$ to enhance the detection power of the resulting classifier. Our numerical experiments demonstrate that this transformation better distinguishes between human- and machine-authored text (see Figure 2). Below, we further provide a simple analytical example to illustrate this power enhancement. (ii) Second, we set the the sampling function $s$ in (2) to the source LLM’s $q$. This allows the numerator to match the form of the partial sum in (3), which enables the application of MCLT for FNR control.
Figure 2: Boxplots visualizing the differences in the statistical measures between human- and LLM-authored passages, comparing AdaDetectGPT (with a learned witness function) and Fast-DetectGPT (without it). A larger positive difference from zero indicates better detection power. As observed, the difference computed by AdaDetectGPT is consistently larger than that of Fast-DetectGPT across the first quartile, median, and third quartile. The left panel shows statistics evaluated on the SQuAD dataset, while the right panel displays results for the WritingPrompts dataset.
An analytical example. Consider a hypothetical “Kingdom of Bit” where all communication uses just two tokens: 1 (yes) and 0 (no). Let $\{p_{t}\}_{t}$ denote the true token distribution of this language. Suppose a malicious wizard creates synthetic citizens who appear similar to ordinary people, but their language follows a simpler distribution $q_{t}(x_{t}|x_{<t})=q(x_{t})$ for some fixed function $q$, being independent of the prior context $x_{<t}$. As we will show later, the detection power of our statistic crucially depends on the following quantity:
$\frac{1}{L}\sum_{t=1}^{L}\big[\mathbb{E}_{\widetilde{X}_{t}\sim q}w(\log q(\widetilde{X}_{t}))-\mathbb{E}_{X_{t}\sim p}w(\log q(X_{t}))\big].$ (5)
The greater the deviation of the expression (5) from zero, the higher the power to detect synthetic humans. When setting $w$ to the identity function, this expression reduces to
$\log\left(\frac{q(1)}{q(0)}\right)\times\left[q(1)-\frac{1}{L}\sum_{t=1}^{L}p_{t}(1)\right],$
where $p_{t}(1)=\mathbb{E}_{X_{<t}\sim p}p(1|X_{<t})$ is determined by the true language distribution $p$. In this case, (5) converges to zero as $q(1)\to 1/2$, regardless of the difference between $q(1)$ and the average of $p_{t}(1)$. As such, there are simple settings in which existing logits-based detectors with an identity witness function will struggle to distinguish between human and machine authored text, independent of the actual distance between $p$ and $q$. However, for any $q(1)\neq 1/2$, there exists a function $w$ that makes (5) equal to $q(1)-L^{-1}\sum_{t=1}^{L}[p_{t}(1)]$, independent of the log ratio (see Appendix B for a formal proof). Thus, whenever $q(1)$ differs from the average $p_{t}(1)$, an appropriate transformation $w$ can reliably detect synthetic humans. We will discuss how to properly learn this witness function in (c).
(b) Classification threshold. Given a witness function $w$, we aim to determine a threshold $c$ so that the FNR of the classifier $T_{w}(\bm{X})>c$ is below a specified level $\alpha>0$. A key observation is that, when the passage $\bm{X}$ is generated from the source LLM, $T_{w}(\bm{X})$ in (4) can be represented by the partial sum in (3) with $Z_{t}=w(\log q_{t}(X_{t}|X_{<t}))$ and $\mathcal{F}_{t}=X_{<t}$. It follows from the MCLT that
$\textrm{FNR}_{w}\coloneqq\mathbb{P}_{\bm{X}\sim q}(T_{w}(\bm{X})\leq c)\to\Phi(c),\quad\textrm{as}~L\to\infty,$ (6)
[…]
We validate this threshold selection both theoretically and empirically. Specifically, Theorem 2 in (e) establishes a finite-sample error bound for our classifier’s FNR. Figure 3 illustrates the effectiveness of FNR control and normal approximation across three benchmark datasets and two language models.
Figure 3: Top panel: FNR of the classifier plotted against the significance level $\alpha$. Bottom panel: Distribution of statistics evaluated on LLM-generated text. The dashed red line is the density function of standard normal random variable. Results are shown across three datasets (from left to right) and two language models (indicated by different colors).
(c) Learning the witness function. With the classifier’s FNR fixed asymptotically at level $\alpha$, one can identify the optimal witness function $w$ by maximizing its TNR, defined by
$\textrm{TNR}_{w}\coloneqq\mathbb{P}_{\bm{X}\sim p}(T_{w}(\bm{X})\leq z_{\alpha}),$ (7)
where the probability is evaluated under the human-generated text distribution $p$. Toward that end, we employ MCLT again to derive a closed-form expression of (7).
By definition, we can represent $T_{w}(\bm{X})$ by the difference $T_{w}^{(1)}(\bm{X})-T_{w}^{(2)}(\bm{X})$ where
$T_{w}^{(1)}(\bm{X})$ $=$ $\frac{\sum_{t=1}^{T}[\log w(X_{t}|X_{<t})-\mathbb{E}_{\widetilde{X}_{t}\sim p_{t}}\log w(\widetilde{X}_{t}|X_{<t})]}{\sqrt{\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(\log w(\widetilde{X}_{t}|X_{<t})))}},$
$T_{w}^{(2)}(\bm{X})$ $=$ $\frac{\sum_{t}[\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))-\mathbb{E}_{\widetilde{X}_{t}\sim p_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))]}{\sqrt{\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}}.$
Here, to ease notations, $p_{t}$ and $q_{t}$ in the expectation and variance are implicitly taken with respect to their conditional distributions $p_{t}(\bullet|X_{<t})$ and $q_{t}(\bullet|X_{<t})$ given $X_{<t}$.
Notice that $T_{w}^{(1)}$ is very similar to $T_{w}$ — the only difference lies in the centering term: the conditional expectation is taken with respect to $p_{t}$ instead of $q_{t}$. Under an equal variance condition where the variance terms in the denominator evaluated at $q_{t}$ and $p_{t}$ converge to the same quantity asymptotically, we can invoke the MCLT to prove the asymptotic normality of $T_{w}^{(1)}(\bm{X})$ under the human distribution $p$. As a result, $\textrm{TNR}_{w}$ can be approximated by
$\mathbb{P}_{\bm{X}\sim p}(T_{w}^{(1)}(\bm{X})\leq z_{\alpha}+T_{w}^{(2)}(\bm{X}))\approx\mathbb{E}_{\bm{X}\sim p}\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X})).$ (8)
Given the external human-written text dataset $\mathcal{H}$, one can estimate the right-hand-side of (8) by $n^{-1}\sum_{i=1}^{n}\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}^{(i)}))$ and optimize this estimator to compute the witness function $\widehat{w}$. However, the resulting $\widehat{w}$ depends on the choice of $\alpha$, which limits its flexibility. To eliminate this dependence, we derive a lower bound on the TNR in the following theorem.
Theorem 1 (TNR lower bound).
Under the equal variance condition and a stochastic dominance condition specified in Appendix D.1, $\textrm{TNR}_{w}$ is asymptotically lower bounded by $\min\{\alpha+\Phi^{\prime}(z_{\alpha})T_{w}^{(2*)},1-\alpha\}$ where $\Phi^{\prime}$ is the derivative of $\Phi$ and $T_{w}^{(2*)}$ denotes a population version of $T_{w}^{(2)}(\bm{X})$, given by
$T_{w}^{(2*)}=\frac{\sum_{t}[\mathbb{E}_{X_{<t}\sim p,\widetilde{X}_{t}\sim q_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))-\mathbb{E}_{X_{<t}\sim p,\widetilde{X}_{t}\sim p_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))]}{\sqrt{\sum_{t}\mathbb{E}_{X_{<t}\sim p}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}}.$ (9)
Compared to $T_{w}^{(2)}(\bm{X})$, both the numerator and denominator of $T_{w}^{(2*)}$ are defined by taking expectations with respect to $X_{<t}\sim p$ for each $t$. In the analytical example, the numerator simplifies to (5) after appropriate scaling. According to Theorem 1, optimizing the lower bound is equivalent to maximizing $T_{w}^{(2*)}$, whose solution is independent of $\alpha$. We also remark that the maximal value $\max_{w}T_{w}^{(2*)}$ is similar to certain integral probability metrics (Müller, 1997) such as the maximum mean discrepancy measure widely studied in machine learning (see e.g., Gretton et al., 2012).
Motivated by Theorem 1, we replace the expectations $\mathbb{E}_{X_{<t}\sim p}$ in both the numerator and denominator of $T_{w}^{(2*)}$ with their empirical average over the dataset $\mathcal{H}$, denote the resulting estimator by $\widehat{T}_{w}^{(2)}$ and compute $\widehat{w}=\arg\max_{w\in\mathcal{W}}\widehat{T}_{w}^{(2)}$ over a function class $\mathcal{W}$. Since $w$ is a one-dimensional function over the space of real-valued logits, the optimization is relatively simple.
Specifically, we adopt a linear function class $\mathcal{W}=\{w(z)=\phi(z)^{\top}\beta:\|\beta\|_{2}=1\}$ for some bounded $d$-dimensional feature mapping $\phi$. In this case, $\widehat{T}_{w}^{(2)}$ simplifies to $\beta^{\top}\psi/\sqrt{\beta^{\top}\Sigma\beta}$ for some $d$-dimensional vector $\psi$ and $d\times d$ semi-definite positive matrix $\Sigma$ (see Appendix C.2 for the detailed derivation). With some calculations, our estimated regression coefficients $\widehat{\beta}$ can be efficiently obtained by solving the linear system $\Sigma\beta=\psi$, leading to $\widehat{w}=\widehat{\beta}^{\top}\phi$. We set $\phi$ to the B-spline basis function (De Boor, 1978) in our implementation and relegate additional details to Appendix C.2.
(d) Extension to black-box settings. When the target LLM’s logits are unavailable, we propose to employ multiple source LLMs with distributions $q_{1},q_{2},\ldots,q_{m}$, rather than a single source model, to enhance the detection power of AdaDetectGPT. Specifically, for each source distribution $q_{i}$, we apply the procedure in (c) to learn a corresponding witness function $\widehat{w}_{i}$, resulting in the statistic $T_{\widehat{w}_{i}}(\bm{X})$. We next take the maximum over all source models and use $\max_{i}T_{\widehat{w}_{i}}(\bm{X})$ for LLM detection. We empirically investigate this procedure in Section 4.
Figure 4: A summary of our theories.
(e) Statistical guarantees. In the following, we establish finite-sample bounds for the FNR, TNR, FPR and TPR of the proposed classifier in the white-box setting; see Figure 4 for a summary of our theories. Recall that $L$ denotes the number of tokens in a passage, $n$ denotes the number of human-authored passages in the training data and $\alpha$ is the target FNR level we wish to control. We also define $w^{*}=\arg\max_{w\in\mathcal{W}}T_{w}^{(2*)}$ as the population limit of our estimated witness function $\widehat{w}$. Finally, let $V_{L}$ denote the ratio
$\frac{L^{-1}\sum_{t=1}^{L}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(\widehat{w}(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}{L^{-1}\sum_{t=1}^{L}\textrm{Var}_{\bm{X}\sim q}(\widehat{w}(\log q_{t}(X_{t}|X_{<t})))}.$
Theorem 2 (FNR).
[…]
Theorem 3 (TNR).
[…]
Corollary 4 (TPR).
Under the condition in Theorem 2, the expected TPR of our classifier is lower bounded by $\mathbb{E}(\textrm{TPR}_{\widehat{w}})=1-\mathbb{E}(\textrm{FNR}_{\widehat{w}})\geq 1-\alpha-O\big(L^{-1/2}\log L\big)-O\big(\mathbb{E}|V_{L}-1|^{1/3}\big)$.
Corollary 5 (FPR).
Under the condition in Theorem 3, the expected FPR of our classifier is upper bounded by $\mathbb{E}(\textrm{FPR}_{\widehat{w}})=1-\mathbb{E}(\textrm{TNR}_{\widehat{w}})\leq 1-\text{TNR}_{w^{*}}+O\big(d^{\gamma}/\sqrt{n}\big)$.
We make a few remarks:
  • Theorem 2 provides an upper bound on the difference between our classifier’s expected FNR and the target level $\alpha$. This excess FNR depends on two factors: (i) the number of tokens $L$, and (ii) the variance ratio $V_{L}$. As $L$ diverges to infinity and $V_{L}$ stabilizes to $1$, the FNR of our classifier approaches the target level. We remark that assumptions similar to the bounded denominator condition in Theorem 2 are commonly imposed (see e.g., Bolthausen, 1982a; Hall & Heyde, 2014).
  • Theorem 3 establishes an upper bound on the difference between our classifier’s TNR and that of the oracle classifier, which has access to the population-level optimal witness function $w^{*}$ . This difference depends on: (i) the training sample size $n$, and (ii) the dimensionality $d$ of the feature mapping $\phi$. Since $\phi$ is defined over the space of one-dimensional log probabilities of tokens, rather than the token space itself, $d$ is independent of the vocabulary size, and can be treated as fixed. Consequently, as the training data size $n$ grows to infinity, our classifier’s TNR converges to that of the oracle classifier. According to Theorem 1, with a sufficiently large $\max\limits_{w\in\mathcal{W}}T_{w}^{(2*)}$, the expected TNR can reach up to $1-\alpha$.
  • Corollaries 4 and 5 follow directly from Theorems 2 and 3, due to the relationships between TPR and FNR, and between FPR and TNR.
Together, these results suggest that our classifier is expected to achieve a high AUC as the significance level $\alpha$ varies, which we will verify empirically in the next section.
4 Experiments
We conduct numerical experiments in both white-box and black-box settings to illustrate the usefulness of AdaDetectGPT. To save space, some implementation details are provided in Appendix F.
Datasets. We consider three widely-used datasets for comparing different detectors, including SQuAD for Wikipedia-style question answering (Rajpurkar et al., 2016), WritingPrompts for story generation (Fan et al., 2018), XSum for news summarization (Narayan et al., 2018), Yelp for crowd-sourced product reviews (Zhang et al., 2015), and Essay for high school and university-level essays (Verma et al., 2024). Following Bao et al. (2024), we randomly sample 500 human-written paragraphs from each dataset and generate an equal number of machine-authored paragraphs by prompting an LLM with the first 120 tokens of the human-written text and requiring it to complete the text with up to 200 tokens. This is a challenging setting where LLM-generated text is mixed with human writing. To evaluate AdaDetectGPT, we train the witness function $\widehat{w}$ on two of the datasets and calculate the AUC on the remaining testing dataset.
Benchmark methods. We compare the proposed AdaDetectGPT against eight state-of-the-art detectors: Likelihood, Entropy, LogRank (Gehrmann et al., 2019), LogRank ratio (LRR, Su et al., 2023), DetectGPT (Mitchell et al., 2023) and its enhanced variants Normalized Perturbed log Rank (NPR, Su et al., 2023) and Fast-DetectGPT (Bao et al., 2024), DNAGPT (Yang et al., 2024b), Binoculars (Hans et al., 2024), RADAR (Hu et al., 2023), ImBD (Chen et al., 2025a), and BiScope (Guo et al., 2024a). We measure the detection power of each detector using AUC.
Results. We first consider the white-box setting where we employ various source models summarized in Table S6 for text generation, with parameter sizes ranging from 1 to 20 billion. Table 1 reports the AUC scores of various detectors across all combinations of datasets and five source models. It can be seen that AdaDetectGPT achieves the highest AUC across all combinations of datasets and source models, outperforming Fast-DetectGPT – the best baseline method – by 12.5%-37%. We also evaluate AdaDetectGPT on three more advanced open-source LLMs: Qwen2.5 (Bai et al., 2025), Mistral (Jiang et al., 2023), and LLaMA3 (Dubey et al., 2024). As shown in Table S7, AdaDetectGPT delivers uniform improvements over Fast-DetectGPT and maintains competitive performance across five datasets, achieving the best results in most cases. These findings highlight the advantage of using an adaptively learned witness function for classification. In Appendix G.2, we further analyze the computational cost of AdaDetectGPT. Training the witness function typically requires less than one minute across different sample sizes and dimensions, with memory usage below 0.5 GB. I In Appendix G.3, we further conduct a sensitivity analysis to investigate the sensitivity of AdaDetectGPT’s AUC score to various factors that may affect the estimation of the witness function. Our results show that AdaDetectGPT is generally robust to the size of training data, the number of B-spline feature mappings, and the distributional shift between training and test data, consistently maintaining superior performance over baselines.
Table 1: AUC scores of various detectors in the white-box setting. The “Relative” rows report the percentage improvement of AdaDetectGPT over Fast-DetectGPT.
Dataset Method Source Model
GPT-2 OPT-2.7 GPT-Neo GPT-J GPT-NeoX Avg.
SQuAD Likelihood 0.7043 0.6872 0.6722 0.6414 0.5948 0.6600
Entropy 0.5513 0.5289 0.5392 0.5385 0.5402 0.5396
LogRank 0.7328 0.7169 0.7047 0.6738 0.6179 0.6892
LRR 0.7703 0.7623 0.7553 0.7286 0.6632 0.5849
NPR 0.8783 0.8342 0.8311 0.7330 0.6454 0.6182
DNAGPT 0.8640 0.8413 0.8106 0.7671 0.6933 0.7953
DetectGPT 0.8565 0.8167 0.8047 0.7124 0.6380 0.6047
Fast-DetectGPT 0.9042 0.8801 0.8731 0.8417 0.7866 0.8571
AdaDetectGPT 0.9265 0.9058 0.9088 0.8618 0.8292 0.8864
Relative ( ) 23.2696 21.4914 28.1306 12.6659 19.9396 20.4909
Writing Likelihood 0.7297 0.7119 0.7032 0.6937 0.6811 0.7039
Entropy 0.5154 0.5132 0.5098 0.5126 0.5275 0.5157
LogRank 0.7510 0.7332 0.7253 0.7141 0.7029 0.7253
LRR 0.7801 0.7599 0.7624 0.7489 0.7362 0.6015
NPR 0.8760 0.8329 0.8373 0.8008 0.7854 0.4942
DNAGPT 0.8962 0.8503 0.8590 0.8450 0.8199 0.8541
DetectGPT 0.8540 0.8060 0.8173 0.7681 0.7514 0.6286
Fast-DetectGPT 0.8972 0.8891 0.8904 0.8879 0.8782 0.8886
AdaDetectGPT 0.9352 0.9175 0.9290 0.9190 0.9158 0.9233
Relative ( ) 36.9089 25.5699 35.2136 27.7142 30.8963 31.1543
XSum Likelihood 0.6410 0.6264 0.6197 0.6086 0.5906 0.6173
Entropy 0.5637 0.5571 0.5864 0.5606 0.5679 0.5671
LogRank 0.6623 0.6493 0.6493 0.6310 0.6109 0.6406
LRR 0.6952 0.6795 0.6967 0.6612 0.6447 0.5361
NPR 0.8211 0.7713 0.8273 0.7676 0.7290 0.6178
DNAGPT 0.7531 0.7283 0.7160 0.6939 0.6634 0.7109
DetectGPT 0.8073 0.7639 0.8244 0.7599 0.7239 0.6110
Fast-DetectGPT 0.8293 0.8067 0.8137 0.7926 0.7622 0.8009
AdaDetectGPT 0.8534 0.8420 0.8532 0.8347 0.8061 0.8379
Relative ( ) 14.1250 18.2659 21.1936 20.3301 18.4492 18.5779
We next consider the black-box setting, where the task is to detect text generated by three widely used advanced LLMs: GPT-4o (Hurst et al., 2024), Claude-3.5-Haiku (Anthropic, 2024), and Gemini-2.5-Flash (Comanici et al., 2025). In this setting, token-level log-probabilities are not publicly accessible. To implement Fast-DetectGPT, we use google/gemma-2-9b and google/gemma-2-9b-it (Team et al., 2024) as the sampling and scoring models, respectively, to construct the classification statistics described in Section 2. The same setup is also applied when implementing AdaDetectGPT. We further benchmark against existing ML-based classifiers, including GPT-2 detectors based on RoBERTa-base and RoBERTa-large provided by OpenAI. The results are reported in Tables 2 and 3. Overall, Fast-DetectGPT remains the strongest baseline, though it is occasionally outperformed by Binoculars and RADAR. AdaDetectGPT consistently improves upon Fast-DetectGPT on several datasets, with gains on XSum, WritingPrompts, and Essay, reaching as high as 37.8% improvement on Essay when detecting Gemini-2.5. Its performance is also comparable on the Yelp dataset. Finally, we evaluate robustness to adversarial attacks in the black-box setting. As shown in Table S10, AdaDetectGPT demonstrates greater resilience than Fast-DetectGPT, showing reduced sensitivity to adversarially perturbed texts.
Table 2: Performance comparison of GPT-4o and Claude-3.5 across datasets.
Method GPT-4o Claude-3.5 XSum Writing Yelp Essay Avg. XSum Writing Yelp Essay Avg. RoBERTaBase 0.5141 0.5352 0.6029 0.5739 0.5655 0.5206 0.5386 0.5630 0.5593 0.5566 RoBERTaLarge 0.5074 0.5827 0.5027 0.6575 0.5626 0.5462 0.6149 0.5105 0.6063 0.5695 Likelihood 0.5194 0.7661 0.8425 0.7849 0.7282 0.6780 0.7502 0.7134 0.6160 0.6894 Entropy 0.5397 0.7021 0.7291 0.6951 0.6665 0.5935 0.6594 0.6625 0.5742 0.6142 LogRank 0.5123 0.7478 0.8259 0.7786 0.7161 0.5109 0.6653 0.7244 0.7111 0.6529 Binoculars 0.9022 0.9572 0.9840 0.9777 0.9552 0.9012 0.9393 0.9752 0.9603 0.9440 RADAR 0.9580 0.8046 0.8558 0.8394 0.8644 0.9187 0.7264 0.8424 0.9152 0.8507 BiScope 0.8333 0.8733 0.9700 0.9600 0.9092 0.8533 0.8800 0.8800 0.9567 0.8925 Fast-DetectGPT 0.9048 0.9588 0.9847 0.9800 0.9571 0.9019 0.9361 0.9768 0.9608 0.9439 AdaDetectGPT 0.9072 0.9611 0.9832 0.9841 0.9589 0.9176 0.9400 0.9728 0.9610 0.9478 Relative ( ) 2.4288 5.6095 20.4444 4.2454 16.0326 6.0543 0.3405 6.9929
Table 3: Performance of Gemini-2.5 across datasets.
Method XSum Writing Yelp Essay Avg. RoBERTaBase 0.5311 0.5202 0.5624 0.7279 0.5854 RoBERTaLarge 0.6583 0.5888 0.6029 0.8180 0.6845 Likelihood 0.7127 0.7547 0.6566 0.7565 0.7201 Entropy 0.5754 0.5088 0.6023 0.6038 0.5726 LogRank 0.6084 0.5743 0.6896 0.7504 0.6557 Binoculars 0.8500 0.9453 0.9698 0.9908 0.9390 RADAR 0.8184 0.5152 0.6300 0.5891 0.6382 BiScope 0.7633 0.6800 0.7097 0.8167 0.7642 Fast-DetectGPT 0.8404 0.9443 0.9695 0.9914 0.9364 AdaDetectGPT 0.8432 0.9484 0.9644 0.9947 0.9377 Relative ( ) 1.7544 7.4163 37.8238 1.9916
Finally, we employ the same five LLMs from the white-box setting to compare Fast-DetectGPT and AdaDetectGPT in settings. Following Bao et al. (2024), we use GPT-J as the source LLM for detecting each of the remaining four target LLMs. Additionally, we implement a variant of AdaDetectGPT that uses GPT-J for scoring and GPT-Neo for sampling. Due to space constraints, the results are presented Table S11 in Appendix G.5, where AdaDetectGPT uniformly outperforms Fast-DetectGPT in all cases, with improvements of up to 29%.
5 Conclusion
We propose AdaDetectGPT, an adaptive LLM detector that learns a witness function $w$ to boost the performance of existing logits-based detectors. A natural approach to learning $w$ is to maximize the TNR of the resulting detector for a fixed FNR level $\alpha$. However, the TNR is a highly complicated function of $\alpha$ and $w$, which makes the learned witness function inherently $\alpha$-dependent — that is, it maximizes the TNR at a particular FNR level but does not guarantee optimality at other FNR levels. To address this, we derive a lower bound on the TNR and propose to learn $w$ by maximizing this lower bound. Notice that the lower bound separates the effects of $\alpha$ and $w$: the witness function $w$ affects the lower bound only through $T_{w}^{(2)*}$, which is independent of $\alpha$. Consequently, the witness function that maximizes this lower bound simultaneously maximizes it across all FNR levels.
In our implementation, we opted to learn the witness function via a B-spline basis due to the straightforward nature of the optimization (finding the optimal witness function boils down to solving a system of linear equations) and its favorable theoretical properties (the estimation error can be shown to attain Stone’s optimal convergence rate, Stone, 1982). The number of -spline basis functions can be selected in a data driven way for instance via Lepski’s method; see Lepski & Spokoiny (1997).
Acknowledgments
Chengchun Shi’s and Jin Zhu’s research was partially supported by the EPSRC grant EP/W014971/1. Hongyi Zhou’s and Ying Yang’s research was partially supported by NSFC 12271286 & 11931001. Hongyi Zhou’s research was also partially supported by the China Scholarship Council.
The authors thank the anonymous referees and the area chair for their insightful and constructive comments, which have led to a significantly improved version of the paper.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We confirm the claims in abstract and introduction are accurately reflected by methodology, theory, and experiments in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitation of the work in Appendix H.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The assumptions are intuitively illustrated in Section 3 and rigorous stated in Appendix D. The proof is attached into Appendix E.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The detailed setting for reproducibility are provided in Section 4 and Appendix F.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The dataset and code in this paper are either publicly available or submitted as a new asset.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The details for experiments are shown in Appendix F.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The results are accompanied by the error bars as can be seen from Figure 2.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provided the compute resources in Appendix FHardware details.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The broader impacts of this work is discussed in Appendix H.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This research does not involve data or models that have a high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: In Section 4 and Appendix F, we have explicitly cited or credited the assets used in the paper and explicitly mentioned the corresponding licenses.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The new asset is the implementation of the methods introduced in the paper. The documentation for the new asset is provided alongside.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The LLM is used only for writing and editing, and it does not impact the core methodology.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Additional related works: Watermarking-based detection
Watermarking embeds subtle signals into LLM-generated text to distinguish it from human-written text (see Ji et al., 2025, Section 4.2, for a recent review of LLM watermarking). Aaronson & Kirchner (2023) propose a watermarking technique based on Gumbel sampling. Follow-up works have focused on preserving text quality during watermarking (Christ et al., 2024; Dathathri et al., 2024; Giboulot & Furon, 2024; Liu & Bu, 2024; Wouters, 2024; Wu et al., 2024), enhancing watermark detection (Dathathri et al., 2024; Huo et al., 2024; Cai et al., 2025) and maintaining robustness against adversarial edits (Golowich & Moitra, 2024).
Our work is related to a line of research that frames watermark detection as a statistical hypothesis testing problem (see e.g., Kirchenbauer et al., 2023; Hu et al., 2024; Kuditipudi et al., 2024; Zhao et al., 2024; Li et al., 2024, 2025; Chen et al., 2025b). Under this framework, rejection of the null hypothesis (that no watermark is present) provides statistical evidence that the text was likely generated by an LLM.
Appendix B Details on the analytic example in Section 3
In this section, we provide rigorous discussion about the analytic example presented in Section 3. Noted that
$\mathbb{E}_{\widetilde{X}_{t}\sim q,X_{<t}\sim p}\left\{w(\log q(X_{t})\right\}$ $=$ $q(1)w(\log q(1))+q(0)w(\log q(0)),$
$\mathbb{E}_{X_{<t+1}\sim p}\left\{w(\log q(X_{t}))\right\}$ $=$ $p_{t}(1)w(\log q(1))+p_{t}(0)w(\log q(0)).$
It follows that
$\mathbb{E}_{\widetilde{X}_{t}\sim q,X_{<t}\sim p}\left\{w(\log q(X_{t})\right\}-\mathbb{E}_{X_{<t+1}\sim p}\left\{w(\log q(X_{t}))\right\}$
$=$ $(q(1)-p_{t}(1))\left[w(\log{q(1)})-w(\log{q(0)})\right].$
If $w$ is an identity function, i.e., $w(x)=x$, then the statistics (5) becomes
$\frac{1}{L}\log\Big(\frac{q(1)}{q(0)}\Big)\sum_{t=1}^{L}(q(1)-p_{t}(1)).$
In this case, (5) converges to zero as $q\to 1/2$ regardless the distribution of $p_{t}$. However, if we consider adaptive witness function, the statistics in (5) becomes
$\frac{1}{L}\left[w(\log q(1))-w(\log q(0))\right]\sum_{t=1}^{L}(q(1)-p_{t}(1)).$
When $q(1)\neq 1/2$ (without generality, we assume q(1)=1q(0)>1/2q(1)=1-q(0)>1/2), there always exists a witness function $w(z)=\mathbb{I}\left\{z>\frac{\log q(1)+\log q(0)}{2}\right\}$ such that (5) becomes
$\frac{1}{L}\left[\mathbb{I}\{\log q(1)>\log q(0)\}-\mathbb{I}\{\log q(0)>\log q(1)\}\right]\sum_{t=1}^{L}(q(1)-p_{t}(1))$
$=$ $\frac{1}{L}\sum_{t=1}^{L}(q(1)-p_{t}(1))=q(1)-\frac{1}{L}\sum_{t=1}^{L}p_{t}(1),$
which is independent of the log ratio.
Appendix C Implementation details
C.1 Data for estimating witness function
In this part, we illustrate how we fetch external human and machined-generate text datasets for training witness function in our experiments.
When testing the performance of AdaDetectGPT on one dataset (e.g., XSum), we recruit the remaining datasets (e.g., SQuAD and WritingPrompt) for training the witness function. This ensures the data for testing would not be included for training. In the white-box setting, since $q^{(s)}$ is known, then the machine-generated text in the training datasets are generated by $q^{(s)}$. In the black-box setting, since $q^{(s)}$ is unknown, the machine-generated text in the training datasets are generated by the surrogate models. In the case where the sampling model and scoring model is different and both using surrogate models, then the generated texts comes from the surrogate scoring model.
C.2 Details: the witness function estimation
In this section we elaborate the estimation of witness function. Recall that the population version of objective function for estimating $w$ is
$\frac{\sum_{t}[\mathbb{E}_{X_{<t}\sim p,\widetilde{X}_{t}\sim q_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))]-\sum_{t}[\mathbb{E}_{X_{<t}\sim p,\widetilde{X}_{t}\sim p_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))]}{\sqrt{\sum_{t}\mathbb{E}_{X_{<t}\sim p}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}},$
and we replace the expectations $\mathbb{E}_{X_{<t}\sim p}$ in both the numerator and denominator with their empirical average over the passages of human. This leads to the following plug-in estimated objective function:
$\frac{\sum_{i}\sum_{t}\mathbb{E}[w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{\sqrt{\sum_{i}\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t})))}}$.
$\approx$ $\frac{\sqrt{2}\sum_{i}\sum_{t}\mathbb{E}[w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t}))]-\sqrt{2}\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{\sqrt{\sum_{i}\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t})))+\sum_{i}\sum_{t}\textrm{Var}(w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t})))}}$,
where $\widetilde{X}_{t}$ is sampled from the machine and the approximation holds because of Assumption 4. Besides, taking the expectation over $\widetilde{X}_{t}$ is time consuming, then we approximate $\mathbb{E}[w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t}))$] and $\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t})))$ with
$\mathbb{E}[w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t}))]\;\textup{ and }\;\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t})))$
where $\widetilde{X}_{<t}$ is sampled from LLM. They can be easily estimated by the Monte Carol method as we can utilize the accessible machine to generate texts. Specifically, we generate a text $\bm{\widetilde{X}}^{(i)}\sim q$. And they are estimated by:
$\mathbb{E}[w(\log q_{t}(\widetilde{X}^{(i)}_{t}|\widetilde{X}^{(i)}_{<t}))]\;\textup{ and }\;\textrm{Var}_{\widetilde{X}^{(i)}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}^{(i)}_{t}|\widetilde{X}^{(i)}_{<t})))$.
After these approximation, we lead to the following two-sample $t$-test type objective that eliminates the constant terms:
$\frac{\sum_{i}\sum_{t}[w(\log q_{t}(\widetilde{X}^{(i)}_{t}|\widetilde{X}^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{\sqrt{\sum_{i}\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}^{(i)}_{<t})))+\sum_{i}\sum_{t}\textrm{Var}(w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t})))}}$
$=$ $\frac{\sum_{i}\sum_{t}[w(\log q_{t}(\widetilde{X}^{(i)}_{t}|\widetilde{X}^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{\sqrt{\sum_{i}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(\sum_{t}w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}^{(i)}_{<t})))+\sum_{i}\textrm{Var}(\sum_{t}w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t})))}}$
For ease notations, we denoted $\log q(X^{(i)}_{t}|X^{(i)}_{<t})$ as $z^{(h)}_{it}$, which is the logit of the $t$-th token of the $i$-th human text. Similarly, for the logit computed from machine-generated text, we define it as $z^{(m)}_{it}$. Using these notations, we get the objective function below:
$\frac{1}{\sqrt{\textup{Var}(\sum_{t}w(z^{(h)}_{t}))+\textup{Var}(\sum_{t}w(z^{(m)}_{t}))}}\left(\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}w(z^{(h)}_{it})-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}w(z^{(m)}_{it})\right)$. (10)
For simplicity, (10) implicitly assumes (i) dataset includes $n$ passages, and (ii) each passage has $L$ tokens. Actually, the computational procedure below can be easily extended to various sample sizes and various tokens number cases.
Recall that the witness function have the form $w(z)=\phi(z)^{\top}\beta$, where $\phi(z)$ are some B-spline basis (De Boor, 1978) and $\beta$ be the parameter. Plug in the linear function to (10), the numerator becomes:
$\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}\beta-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}\beta$
As for the denominator, we estimate $\textup{Var}(\sum_{t}w(z^{(h)}_{t}))$ by $\beta^{\top}\widehat{\Sigma}^{(h)}\beta$, and estimate $\textup{Var}(\sum_{t}w(z^{(m)}_{t}))$ by $\beta^{\top}\widehat{\Sigma}^{(m)}\beta$, where in particular $\widehat{\Sigma}^{(h)}=\sum_{i=1}^{n}\widehat{\Sigma}_{i}^{(h)}$ with
$\widehat{\Sigma}_{i}^{(h)}=\frac{1}{L}(\mathbf{Z}^{(h)}_{i})^{\top}\mathbf{Z}^{(h)}_{i}-\widehat{\mu}^{(h)}_{i}(\widehat{\mu}^{(h)}_{i})^{\top}$,
$\mathbf{Z}_{i}^{(h)}=\left(\phi(z_{i1}^{(h)}),\ldots,\phi(z_{iL}^{(h)})\right)^{\top},$
$\widehat{\mu}_{i}^{(h)}=\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top},$
and $\widehat{\Sigma}^{(m)}=\sum_{i=1}^{n}\widehat{\Sigma}_{i}^{(m)}$; the expression for the $\widehat{\Sigma}_{i}^{(m)}$’s is analogous to that of the $\widehat{\Sigma}_{i}^{(h)}$’s and is therefore omitted. Consequently, the objective function can be rewritten as:
$\frac{\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}\right]\beta}{\sqrt{\beta^{\top}(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})\beta}}$
$=$ $\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}\right]\beta\times\frac{1}{\|(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{1/2}\beta\|_{2}}$
$=$ $\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}\right]\times(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-\frac{1}{2}}\alpha,$
where $\alpha=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{1/2}\beta/\|(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{1/2}\beta\|_{2}$ and thus $\|\alpha\|_{2}$. The maximizer of the objective function has a closed-form solution:
$\widehat{\alpha}=\frac{\widetilde{\alpha}}{\|\widetilde{\alpha}\|_{2}}$
with
$\widetilde{\alpha}=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-\frac{1}{2}}\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})\right];$ (11)
and thus $\|\widehat{\alpha}\|_{2}=1$. For $\alpha$ satisfying $\|\alpha\|_{2}=1$, we have $\beta=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-1/2}\alpha$. Therefore, the expression of $\widehat{\beta}$ is:
$\widehat{\beta}$ $=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-1/2}\widehat{\alpha}$. (12)
and $\widehat{w}(z)=\phi(z)^{\top}\widehat{\beta}$.
From Section G.3, considering the increase of n_base would raise higher computational cost, and thus, we recommend to set n_base=16 and order=2.
Appendix D Assumptions for theories
In this section, we list the assumptions required for the theorems presented in Section 3 to hold, and discuss when they can reasonably be expected to hold in practice and how they may be relaxed.
D.1 Assumptions
We work under the following assumptions on the data generating process.
Assumption 1 (Margin).
With $T_{w}(\bullet)$ defined as in (4) and $w^{*}(\bullet)$ defined as the optimizer of (9), for any $\alpha\in(0,1)$ there are constants $\delta_{\alpha},C_{\alpha}$ depending only on $\alpha$ such that for any $x\leq\delta_{\alpha}$ it holds that $\mathbb{P}_{\boldsymbol{X}\sim p}(|T_{w^{*}}(\boldsymbol{X})-z_{\alpha}|\leq x)\leq C_{\alpha}x$.
We also require the following technical conditions hold in order to obtain TNR lower bound and FNR control (Theorem 1 and Theorem 2).
Assumption 2 (Minimum eigenvalue).
For each $t=1,\dots,L$ introduce the quantities
$\mu_{t}^{(1)}=\mathbb{E}_{X_{<t}\sim p}\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}\phi(\log q_{t}(\widetilde{X}_{t}\mid X_{<t})),$
$\Sigma_{t}=\mathbb{E}_{X_{<t}\sim p}\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}\phi(\log q_{t}(\widetilde{X}_{t}\mid X_{<t}))\phi(\log q_{t}(\widetilde{X}_{t}\mid X_{<t}))^{\top}-\mu_{t}^{(1)}(\mu^{(1)}_{t})^{\top}$.
There are absolute constant $C>0$ and $\gamma>0$ such that $\lambda_{\text{min}}(\Sigma_{t})\geq Cd^{-\gamma}$ for all $t$.
Assumption 3 (Stochastic dominance).
For any witness function $w\in\mathcal{W}$,
$\sum_{t}\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))-\mathbb{E}_{\widetilde{X}_{t}\sim p_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))\geq 0$
holds almost surely.
Assumption 4 (Equal variance).
For any non-constant witness function $w$, define
$\sigma_{q,L}^{2}\coloneqq\frac{1}{L}\sum_{t=1}^{L}\textup{Var}_{\widetilde{X}_{t}\sim q_{t}}\left(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t}))\right),$
$\sigma_{p,L}^{2}\coloneqq\frac{1}{L}\sum_{t=1}^{L}\textup{Var}_{\widetilde{X}_{t}\sim p_{t}}\left(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t}))\right).$
$\sigma_{q,L}^{2},\sigma_{p,L}^{2}$ are lower bounded by some constant $\sigma_{w}^{2}>0$ almost surely. Moreover, $\sigma_{q,L}-\sigma_{p,L}\to 0$ in probability as $L\to\infty$.
Assumption 5.
For any witness function $w$, define
$\bar{\sigma}_{q,L}^{2}=\frac{1}{L}\sum_{t=1}^{L}\textup{Var}_{\bm{X}\sim q}\left(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t}))\right),$
$\bar{\sigma}_{p,L}^{2}=\frac{1}{L}\sum_{t=1}^{L}\textup{Var}_{\bm{X}\sim p}\left(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t}))\right).$
If $\bm{X}\sim q$, then $\bar{\sigma}_{q,L}^{2}/\sigma_{q,L}^{2}\to 1$ in probability. If $\bm{X}\sim p$, then $\bar{\sigma}_{p,L}^{2}/\sigma_{p,L}^{2}\to 1$ in probability.
D.2 Discussion on assumptions
Assumption 1 is commonly assumed in the the classification and inference literature; see, for instance, Audibert & Tsybakov (2007); Shi et al. (2020). For Assumption 3, the witness function is trained to maximize the mean discrepancy. When $w(x)\equiv 1$, the mean discrepancy is exactly zero. Therefore, it is reasonable to restrict the function class to those functions that yield a non-negative mean discrepancy. Assumption 4 basically requires the conditional variance of logits be asymptotically equivalent for human-authored and machine-generated passages. This assumption is not overly restrictive, as the variance discrepancy between the two types of passages is relatively small in our dataset (see Figure 2). Assumption 5 is commonly assumed in martingale central limit theorem literature, see e.g. Hall & Heyde (2014); Bolthausen (1982b). Empirical results further support the validity of Assumption 5: the sample mean of the ratio remains nearly constant as a function of $L$, while its variance (in parentheses) approaches zero as $L\to\infty$ across all three datasets (see Table S4), suggesting that this condition is practical. Furthermore, two standard hypothesis tests — the Kolmogorov–Smirnov (KS) test and the Shapiro–Wilk (SW) test — are conducted to evaluate whether the proposed statistic follows a normal distribution. As shown in Table S5, almost all $p$-values exceed 0.1 (most by a large margin), indicating that our test statistic passes normality tests in most cases. These test results also provide strong empirical support for the validity of MCLT regularity conditions.
$L$ 100 150 200 250 300 350
XSum 1.09(0.12) 1.07(0.09) 1.06(0.08) 1.04(0.07) 1.01(0.06) 0.99(0.05)
SQuAD 1.03(0.10) 1.02(0.07) 1.02(0.06) 1.02(0.05) 1.01(0.05) 1.03(0.04)
Writing 1.10(0.06) 1.09(0.05) 1.09(0.04) 1.08(0.03) 1.07(0.03) 1.05(0.03)
Table S4: Sample mean and variance (in parentheses) of the ratio evaluated on 3 datasets as $L$ increases.
LLM Test XSum SQuAD Writing
GPT-Neo KS 0.72 0.54 0.18
GPT-Neo SW 0.50 0.65 0.89
GPT-2 KS 0.10 0.52 0.28
GPT-2 SW 0.37 0.026 0.14
Table S5: $p$-values of KS and SW tests across 3 datasets and 2 source LLMs.
Appendix E Proofs
E.1 Notations
Throughout the proofs we will make use of the following notation. We will denote absolute constants by $\kappa_{1},\kappa_{2},\cdots$. For a sequence of random variables $\{X_{n}\mid n\geq 1\}$ with distribution functions $\{F_{X_{n}}|n\geq 1\}$ and some (possibly degenerate) random variable $Y$ with distribution function $F_{Y}$ we write $X_{n}\stackrel{{\scriptstyle p}}{{\rightarrow}}Y$ as $n\rightarrow\infty$ if $\lim\limits_{n\rightarrow\infty}\mathbb{P}(|X_{n}-Y|>\delta)=0$ for all $\delta>0$, and $X_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}Z$ if $\lim\limits_{n\rightarrow\infty}F_{X_{n}}(x)=F_{Y}(x)$ at every continuity point of $F_{Y}(\cdot)$. For a vector $x=(x_{1},\dots,x_{d})^{\top}\in\mathbb{R}^{d}$ we write $\|x\|_{p}=(\sum_{j=1}^{d}x_{j}^{p})^{1/p}$ with $0<p<\infty$ for its $\ell_{p}$-norm.
E.2 Preparatory results
Theorem S6 (Bounded differences inequality).
Let $\mathcal{X}$ be a measurable space. A function $f:\mathcal{X}^{n}\to\mathbb{R}$ has the bounded difference property for some constants $c_{1},\dots,c_{n}$, i.e., for each $i=1,\dots,n$,
$\sup_{\begin{subarray}{c}x_{1},\dots,x_{n}\\ x_{i}^{\prime}\in\mathcal{X}\end{subarray}}\left|f\left(x_{1},\dots,x_{i-1},x_{i},x_{i+1},x_{n}\right)-f\left(x_{1},\dots,x_{i-1},x_{i}^{\prime},x_{i+1},\dots x_{n}\right)\right|\leq c_{i}.$ (13)
Then, if $X_{1},\dots,X_{n}$ is a sequence of identically distributed random variables and (13) holds, putting $Z=f\left(X_{1},\dots,X_{n}\right)$ and $\nu=\frac{1}{4}\sum_{i=1}^{n}c_{i}^{2}$ for any $t>0$, it holds that
$\mathbb{P}\left(Z-\mathbb{E}\left(Z\right)>t\right)\leq e^{-t^{2}/\left(2\nu\right)}.$
Proof of Theorem S6.
See Section 2 in Wainwright (2019). ∎
Theorem S7 (Martingale central limit theorem).
Let $\left\{M_{n,i}\mid 1\leq i\leq k_{n},n\geq 1\right\}$ be a zero mean square integrable martingale array with respect to the filtrations $\left\{\mathcal{F}_{n,i}\mid 1\leq i\leq k_{n},n\geq 1\right\}$ having increments $X_{n,i}=M_{n,i}-M_{n,i-1}$. If the following conditions hold
  • C1: $\sum_{i=1}^{k_{n}}\mathbb{E}\left[X_{n,i}\mathbf{1}_{\left\{\left|X_{n,i}\right|>\delta\right\}}\mid\mathcal{F}_{n,i-1}\right]\stackrel{{\scriptstyle p}}{{\rightarrow}}0$ as $n\rightarrow\infty$ for all $\delta>0$
  • C2: $\sum_{i=1}^{k_{n}}\mathbb{E}\left[X_{n,i}^{2}\mid\mathcal{F}_{n,i-1}\right]\stackrel{{\scriptstyle p}}{{\rightarrow}}\sigma^{2}$ as $n\rightarrow\infty$
  • C3: the $\sigma$-fields are nested: $\mathcal{F}_{n,i}\subseteq\mathcal{F}_{n+1,i}$ for $1\leq i\leq k_{n}$ and $n\geq 1$
then $M_{n,k_{n}}\stackrel{{\scriptstyle d}}{{\rightarrow}}Z$ as $n\rightarrow\infty$, where $Z\sim\mathcal{N}\left(0,\sigma^{2}\right)$.
Proof.
See Corollary 3.1 in Hall & Heyde (2014) and Theorem 2 in Bolthausen (1982a). ∎
Lemma S1 (Convergence rate of martingale central limit theorem).
Let $\bm{X}=(X_{1},\ldots X_{n})$ be sequences of real valued random variables satisfying for all $1\leq t\leq n$
$\mathbb{E}\left\{X_{t}|X_{<t}\right\}=0\quad\textup{almost surely.}$
Let $\sigma_{t}^{2}=\mathbb{E}\left\{X_{t}^{2}\big|X_{<t}\right\}$ and $\bar{\sigma}_{t}^{2}=\mathbb{E}\left\{X_{t}^{2}\right\}$. Denote $s_{n}^{2}=\sum_{t=1}^{n}\bar{\sigma}_{t}^{2}$, $V_{n}^{2}=\sum_{t=1}^{n}\sigma_{t}^{2}/s_{n}^{2}$. Suppose $\|X_{n}\leq\gamma$ almost surely for all $n$, $s_{n}^{2}/n\to s^{2}$, and $V_{n}^{2}\to 1$ as $n\to\infty$, then
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\leq z\right)-\Phi(z)\right|\leq O\left(\frac{\log n}{\sqrt{n}s^{3}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right),$
where $\Phi(\bullet)$ is the cumulative distribution function of standard normal distribution.
Proof.
From Corollary 1 in Bolthausen (1982a), we have
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{s_{n}}\leq z\right)-\Phi(z)\right|\leq O\left(\frac{\log n}{\sqrt{n}s^{3}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right).$
It follows that
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\leq z\right)-\Phi(z)\right|$ (14)
$=$ $\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{s_{n}}\leq z+\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\right)(V_{n}-1)\right)-\Phi(z)\right|$
$\leq$ $\sup_{z\in\mathbb{R}}\mathbb{E}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{s_{n}}\leq z+\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\right)(V_{n}-1)\right)-\Phi\left(z+\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\right)(V_{n}-1)\right)\right|$
$+\sup_{z\in\mathbb{R}}\mathbb{E}\left|\Phi\left(z+\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\right)(V_{n}-1)\right)-\Phi(z)\right|$
$\leq$ $O\left(\frac{\log n}{\sqrt{n}s^{3}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right)+\sup_{z\in\mathbb{R}}|\nabla\Phi(z)|\times\mathbb{E}\left|\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\right)(V_{n}-1)\right|.$
Using the definition $\Phi(z)$, we know that
$\sup_{z\in\mathbb{R}}|\nabla\Phi(z)|=\sup_{z\in\mathbb{R}}\frac{1}{\sqrt{2\pi}}\exp(-z^{2}/2)\leq\frac{1}{\sqrt{2\pi}},$ (15)
and using the fact that $\|X_{t}\|_{\infty}\leq\gamma$, $s_{n}^{2}/n\to s^{2}>0$ and $V_{n}\to 1$ in probability, we obtain that
$\mathbb{E}\left|\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\right)(V_{n}-1)\right|$ $=$ $\mathbb{E}\left|\left(\frac{\sum_{t=1}^{n}X_{t}}{s_{n}}\right)V_{n}(V_{n}-1)\right|$ (16)
$\leq$ $\left(\mathbb{E}\left\{\frac{\left(\sum_{t=1}^{n}X_{t}\right)^{2}}{s_{n}^{2}}V_{n}^{2}\right\}\mathbb{E}\left\{(V_{n}-1)^{2}\right\}\right)^{1/2}$
$=$ $O((\mathbb{E}|V_{n}^{2}-1|)^{1/2}),$
where the inequality is Cauchy inequality. Combining equations (14), (15) and (16), we obtain
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\leq z\right)-\Phi(z)\right|\leq O\left(\frac{\log n}{\sqrt{n}s^{3}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right),$ (17)
which finishes the proof. ∎
Lemma S2.
Suppose $X$ is a random variable. Let $\Phi(\bullet)$ be the cumulative distribution function of standard normal distribution and $\Phi^{\prime}(\bullet)$ be its derivative. Then for any random variable $X$,
$\mathbb{E}\Phi(z_{\alpha}+X)\geq\min\{1-\alpha,\alpha+\Phi^{\prime}(z_{\alpha})\mathbb{E}X\},$
where $0<\alpha<1/2$, $z_{\alpha}$ is the $\alpha$-th quantile of standard normal distribution.
Proof of Lemma S2.
Since $\phi(x)=(\sqrt{2\pi})^{-1}\exp\left(-x^{2}/2\right)$, we noted that $\phi(x)\geq\phi(z_{\alpha})$ holds if and only if $z_{\alpha}\leq x\leq z_{1-\alpha}$. Therefore, if $0\leq X<z_{1-\alpha}-z_{\alpha}$, then by the mean value theorem,
$\Phi(z_{\alpha}+X)$ $=$ $\Phi(z_{\alpha})+\Phi^{\prime}(\xi)X\geq\alpha+\Phi^{\prime}(z_{\alpha})X,$
where $\xi$ lies between $z_{\alpha}$ and $z_{1-\alpha}$. If $X\leq 0$, then
$\Phi(z_{\alpha}+X)$ $=$ $\Phi(z_{\alpha})+\Phi^{\prime}(\eta)X\geq\alpha+\Phi^{\prime}(z_{\alpha})X,$
where $\eta$ lies between $X$ and $z_{\alpha}$. Moreover, if $X\geq z_{1-\alpha}-z_{\alpha}$, then $z_{\alpha}+X\geq z_{1-\alpha}$, It follows that $\Phi(z_{\alpha}+X)\geq\Phi(z_{1-\alpha})=1-\alpha$. Therefore,
$\mathbb{E}\Phi(z_{\alpha}+X)$ $\geq$ $\mathbb{E}\min\left\{\alpha+\Phi^{\prime}(z_{\alpha})X,1-\alpha\right\}$
$\geq$ $\min\left\{\alpha+\Phi^{\prime}(z_{\alpha})\mathbb{E}X,1-\alpha\right\},$
where the last inequality follows from Jensen’s inequality. This finishes the proof. ∎
The following Lemma is essential to derive a sharper lower bound for Theorem 1.
Lemma S3.
Suppose $X$ is a random variable which satisfies $X\geq 0$ almost surely. Let $\Phi(\bullet)$ be the cumulative distribution function of standard normal distribution and $\phi(\bullet)$ be the density function of standard normal distribution. Then for any random variable $X$,
$\mathbb{E}\Phi(z_{\alpha}+X)\geq\sup_{\beta\in(0,\alpha]}\min\{1-\beta,\alpha+\Phi^{\prime}(z_{\beta})\mathbb{E}X\},$
where $0<\alpha<1/2$, $z_{\alpha}$ is the $\alpha$-th quantile of standard normal distribution.
Proof of Lemma S3.
Since $\Phi^{\prime}(x)=(\sqrt{2\pi})^{-1}\exp\left(-x^{2}/2\right)$, we noted that if $0<\alpha<1/2$, then for any $\beta\in(0,\alpha]$, $\Phi^{\prime}(x)\geq\Phi^{\prime}(z_{\beta})$ holds if and only if $z_{\beta}\leq x\leq z_{1-\beta}$. Therefore, if $0\leq X<z_{1-\beta}-z_{\alpha}$, then by mean value theorem,
$\Phi(z_{\alpha}+X)$ $=$ $\Phi(z_{\alpha})+\Phi^{\prime}(\xi)X\geq\alpha+\Phi^{\prime}(z_{\beta})X,$
where $\xi$ lies between $z_{\alpha}$ and $z_{1-\beta}$. Moreover, if $X\geq z_{1-\beta}-z_{\alpha}$, then $z_{\alpha}+X\geq z_{1-\beta}$, It follows that $\Phi(z_{\alpha}+X)\geq\Phi(z_{1-\beta})=1-\beta$. Therefore,
$\mathbb{E}\Phi(z_{\alpha}+X)$ $\geq$ $\mathbb{E}\min\left\{\alpha+\Phi^{\prime}(z_{\beta})X,1-\beta\right\}$
$\geq$ $\min\left\{\alpha+\Phi^{\prime}(z_{\beta})\mathbb{E}X,1-\beta\right\},$
where the last inequality follows from Jensen’s inequality. Take supremum with respect to $\beta\in(0,\alpha]$ on both sides, we obtain
$\mathbb{E}\Phi(z_{\alpha}+X)\geq\sup_{\beta\in(0,\alpha]}\min\{1-\beta,\alpha+\Phi^{\prime}(z_{\beta})\mathbb{E}X\}.$
This finishes the proof. Noted that by taking $\beta=\alpha$, the conclusion corresponds to Lemma S2. ∎
In Lemma S4 below, we provide a bound for the parameter estimation. Before doing so, we introduce some related functions
$Q_{*}\left(\beta\right)=\left\{\beta^{\top}\Sigma\beta\right\}^{-\frac{1}{2}}\beta^{\top}\mu$ (18)
$\widehat{Q}_{n}\left(\beta\right)=\left\{\beta^{\top}\widehat{\Sigma}_{n}\beta\right\}^{-\frac{1}{2}}\beta^{\top}\widehat{\mu}_{n},\hskip 10.00002ptn\in\mathbb{N}$ (19)
where in particular we have put $\widehat{\mu}_{n}=L^{-1}\sum_{t=1}^{L}\widehat{\mu}_{t}^{(1)}-\widehat{\mu}_{t}^{(2)}$ and $\mu=L^{-1}\sum_{t=1}^{L}\mu^{(1)}_{t}-\mu^{(2)}_{t}$ with
$\widehat{\mu}_{t}^{(1)}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}\phi\left(\log q_{t}\left(\widetilde{X}_{t}\mid X_{<t}^{(i)}\right)\right)\hskip 10.00002pt$
$\widehat{\mu}_{t}^{(2)}=\frac{1}{n}\sum_{i=1}^{n}\phi\left(\log q_{t}\left(X_{t}^{(i)}\mid X_{<t}^{(i)}\right)\right)$
$\mu_{t}^{(1)}=\mathbb{E}_{X_{<t}\sim p}\mathbb{E}_{X_{t}\sim q_{t}}\phi\left(\log q_{t}\left(X_{t}\mid X_{<t}\right)\right)$
$\mu_{t}^{(2)}=\mathbb{E}_{X_{<t}\sim p}\mathbb{E}_{X_{t}\sim p_{t}}\phi\left(\log q_{t}\left(X_{t}\mid X_{<t}\right)\right)$
for each $t=1,\dots,L$, as well as $\widehat{\Sigma}_{n}=L^{-1}\sum_{t=1}^{L}\widehat{\Sigma}_{t}$ and $\Sigma=L^{-1}\sum_{t=1}^{L}\Sigma_{t}$ with
$\widehat{\Sigma}_{t}=\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}\left[\phi\left(\log q_{t}\left(\widetilde{X}_{t}\mid X_{<t}^{(i)}\right)\right)\phi\left(\log q_{t}\left(\widetilde{X}_{t}\mid X_{<t}^{(i)}\right)\right)^{\top}\right]-\widehat{\mu}_{t}^{(1)}\left(\widehat{\mu}_{t}^{(1)}\right)^{\top}$
for each $t=1,\dots,L$.
Lemma S4.
Grant the assumptions in Section D hold. Let $\beta^{*}$ be the maximizer of the function (18) over all $\beta$’s with $\ell_{2}$ norm equal to $1$ and let $\widehat{\beta}$ be the maximizer of the empirical counterpart (19). There are absolute constants $\kappa_{1}$ and $\kappa_{2}$ depending only on the constants stated in the assumptions such that for any $z>0$ it holds that $\|\widehat{\beta}-\beta^{*}\|_{2}\leq z$ with probability at least $1-\kappa_{1}\exp\left(-\kappa_{2}d^{-5\gamma}n(\min\{z,1\})^{2}\right)$.
Proof.
Observing the $\widehat{\beta}\in\operatorname*{arg\,max}\limits_{\beta}\widehat{Q}_{n}(\beta)$ and $\beta^{*}\in\operatorname*{arg\,max}\limits_{\beta}Q^{*}(\beta)$ for any $z>0$
$\mathbb{P}\left(\left\|\widehat{\beta}-\beta^{*}\right\|_{2}>z\right)=$ $\mathbb{P}\left(\sup_{\beta:\left\|\beta-\beta^{*}\right\|_{2}>z}Q_{n}\left(\beta\right)-Q_{n}\left(\beta^{*}\right)>0\right)$ (20)
$\leq$ $\mathbb{P}\left(\sup_{\beta:\left\|\beta-\beta^{*}\right\|_{2}>z}\left|Q_{n}\left(\beta\right)-Q_{*}\left(\beta\right)\right|>\frac{1}{2}\inf_{\beta:\left\|\beta-\beta^{*}\right\|_{2}>z}\left[Q_{*}\left(\beta^{*}\right)-Q_{*}\left(\beta\right)\right]\right)$
$+\mathbb{P}\left(\left|Q_{n}\left(\beta^{*}\right)-Q_{*}\left(\beta^{*}\right)\right|>\frac{1}{2}\inf_{\beta:\left\|\beta-\beta^{*}\right\|_{2}>z}\left[Q_{*}\left(\beta^{*}\right)-Q_{*}\left(\beta\right)\right]\right).$
For $\beta$ such that $\left\|\beta-\beta^{*}\right\|_{2}>z$, due to Assumption 2 with some absolute $\kappa_{1}$ we must have that
$Q_{*}\left(\beta^{*}\right)-Q_{*}\left(\beta\right)$ $\geq\kappa_{1}d^{-\gamma}\left\|\beta-\beta^{*}\right\|_{2}\geq\kappa_{1}zd^{-\gamma}.$ (21)
For any $\beta\in\mathbb{R}^{d}$ with $\left\|\beta\right\|_{2}=1$ it holds that
$\left|Q_{n}\left(\beta\right)-Q_{*}\left(\beta\right)\right|$
$\leq$ $\left(\beta^{\top}\Sigma\beta\right)^{-\frac{1}{2}}\left|\beta^{\top}\left(\widehat{\mu}_{n}-\mu\right)\right|+\left|\beta^{\top}\widehat{\mu}_{n}\right|\left|\left(\beta^{\top}\widehat{\Sigma}_{n}\beta\right)^{-\frac{1}{2}}-\left(\beta^{\top}\Sigma\beta\right)^{-\frac{1}{2}}\right|$
$\leq$ $\left(\beta^{\top}\Sigma\beta\right)^{-\frac{1}{2}}\left|\beta^{\top}\left(\widehat{\mu}_{n}-\mu\right)\right|$
$\hskip 10.00002pt+\frac{1}{2}\left\{\min\left(\beta^{\top}\widehat{\Sigma}_{n}\beta,\beta^{\top}\Sigma\beta\right)\right\}^{-\frac{3}{2}}\left|\beta^{\top}\widehat{\mu}_{n}\right|\left|\beta^{\top}\widehat{\Sigma}_{n}\beta-\beta^{\top}\Sigma\beta\right|$ (22a)
$\leq$ $\left\{\lambda_{\text{min}}(\Sigma)\right\}^{-\frac{1}{2}}\left|\beta^{\top}\left(\widehat{\mu}_{n}-\mu\right)\right|$ (22b)
$\hskip 10.00002pt+\frac{1}{2}\left\{\min\left(\lambda_{\text{min}}(\widehat{\Sigma}_{n}),\lambda_{\text{min}}(\Sigma)\right)\right\}^{-\frac{3}{2}}\left|\beta^{\top}\widehat{\mu}_{n}\right|\left|\beta^{\top}\widehat{\Sigma}_{n}\beta-\beta^{\top}\Sigma\beta\right|,$ (22c)
where in particular (22a) holds due to the inequality
$\left|\frac{1}{\sqrt{x}}-\frac{1}{\sqrt{y}}\right|=\left|\frac{x-y}{\sqrt{xy}(\sqrt{x}+\sqrt{y})}\right|\leq\frac{1}{2\left(\min(x,y)\right)^{\frac{3}{2}}}\left|x-y\right|$ (23)
for $x,y>0$. For (22b) note that due to the boundedness of $\phi(\bullet)$ for any $\beta$ with $\|\beta\|_{2}=1$ it holds that $\beta^{\top}\left(\widehat{\mu}_{n}-\mu\right)$ is the average of $n$ i.i.d. random variables each bounded in absolute value by a constant which does not depend on $\beta$. Therefore lower bounding $\lambda_{\text{min}}(\Sigma)\geq\kappa_{2}d^{-\gamma}$ for some absolute $\kappa_{2}$ using again the boundedness of $\phi(\bullet)$ and applying Hoeffding’s inequality, for any $z>0$
[…]
for certain absolute $\kappa_{3},\kappa_{4}$. The following argument will be valid on the event
$\lambda_{\text{min}}(\widehat{\Sigma}_{n})\geq\frac{1}{2}\lambda_{\text{min}}(\Sigma).$ (25)
For (22c) notice first that with $\|\beta\|_{2}=1$ the quantity $|\beta^{\top}\widehat{\mu}_{n}|$ is almost surely bounded from above by some absolute constant independent of $\beta$ and $d$. Moreover due to the boundedness of $\phi(\bullet)$ it is easy to see that the statistic $|\beta^{\top}\widehat{\Sigma}_{n}\beta-\beta^{\top}\Sigma\beta|$ is a self-bounding function of $n$ random variables with constants (see equation (13)) $c_{i}\propto\frac{1}{n}$ for $i=1,\dots,n$ which again do not depend on $\beta$. Therefore, on the event (25) applying Theorem S6 we obtain that
[…]
for certain absolute $\kappa_{5},\kappa_{6}$. Finally, note that
$\lambda_{\text{min}}(\widehat{\Sigma})=\min_{\beta:\|\beta\|_{2}=1}\beta^{\top}\widehat{\Sigma}_{n}\beta\geq\lambda_{\text{min}}(\Sigma)-\max_{\beta:\|\beta\|_{2}=1}\beta^{\top}(\widehat{\Sigma}_{n}-\Sigma)\beta$ (27)
and arguing as in (26), the final term (27) is no larger than $\frac{1}{2}\kappa_{2}d^{-\gamma}$ with probability at least
$1-2\exp\left(-\frac{\kappa_{7}n}{d^{2\gamma}}\right).$
Since by Assumption 2 we must have that $\kappa_{2}d^{-\gamma}\leq\lambda_{\text{min}}(\Sigma)$ the event (25) must hold with the above probability. Since the above arguments hold for any $\beta$ with $\|\beta\|_{2}=1$, plugging (24) and (26) back into (20) and accounting for the event (27) the stated result follows. ∎
E.3 Proof of Theorem 1
According to the decomposition $T_{w}(\bm{X})=T_{w}^{(1)}(\bm{X})-T_{w}^{(2)}(\bm{X})$ with $T_{w}^{(1)}(\bm{X}),T_{w}^{(2)}(\bm{X})$ defined by
$T_{w}^{(1)}(\bm{X})$ $=$ $\frac{\sum_{t}[w(\log q_{t}(X_{t}|X_{<t}))-\mathbb{E}_{\widetilde{X}_{t}\sim p_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))]}{\sqrt{\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}}$
$T_{w}^{(2)}(\bm{X})$ $=$ $\frac{\sum_{t}[\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))-\mathbb{E}_{\widetilde{X}_{t}\sim p_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))]}{\sqrt{\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}},$ (28)
we obtain that the TNR can be represented as
$\mathbb{P}_{\bm{X}\sim p}\left(T_{w}(\bm{X})\leq z_{\alpha}\right)$ $=$ $\mathbb{P}_{\bm{X}\sim p}\left(T_{w}^{(1)}(\bm{X})\leq z_{\alpha}+T_{w}^{(2)}(\bm{X})\right)$ (29)
It is easy to verify that when $\bm{X}\sim p$, $T_{w}^{(1)}(\bm{X})\sigma_{q,L}/\sigma_{p,L}$ converges to standard normal distribution, using the convergence rate for martingale central limit theorem (i.e., Theorem S1), we obtain that
$\mathbb{P}_{\bm{X}\sim p}\left(T_{w}(\bm{X})\leq z_{\alpha}\right)$ $=$ $\mathbb{P}_{\bm{X}\sim p}\left(T_{w}^{(1)}(\bm{X})\frac{\sigma_{q,L}}{\sigma_{p,L}}\leq(z_{\alpha}+T_{w}^{(2)}(\bm{X}))\frac{\sigma_{q,L}}{\sigma_{p,L}}\right)$
$\geq$ $\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}))+\left(\Phi\left((z_{\alpha}+T_{w}^{(2)}(\bm{X}))\frac{\sigma_{q,L}}{\sigma_{p,L}}\right)-\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}))\right)$
$\qquad+O\left(\log L/\sqrt{L}\right)$
$\geq$ $\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}))-\sup_{z\in\mathbb{R}}\Phi^{\prime}(z)\times\left|z_{\alpha}+T_{w}^{(2)}(\bm{X})\right|\times\left|\frac{\sigma_{q,L}}{\sigma_{p,L}}-1\right|$
$\qquad+O\left(\log L/\sqrt{L}\right).$
Take expectation on both sides, we obtain
$\mathbb{P}_{\bm{X}\sim p}\left(T_{w}(\bm{X})\leq z_{\alpha}\right)\geq\mathbb{E}\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}))+o(1)+O(\log L/\sqrt{L}),$
because of Assumption 4. Now define $\widetilde{\sigma}_{q,L}^{2}=\mathbb{E}_{\bm{X}\sim p}\sigma_{q,L}^{2}$. Then under equal variance assumption (Assumption 4), we also have $\sigma_{q,L}-\widetilde{\sigma}_{q,L}\to 0$ in probability. It follows that for any $\epsilon>0$,
$\mathbb{E}\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}))$ $=$ $\mathbb{E}\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}))\mathbb{I}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|\leq\epsilon\}$
$\qquad+\mathbb{E}\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}))\mathbb{I}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|>\epsilon\}$
$\geq$ $\mathbb{E}\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}))\mathbb{I}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|\leq\epsilon\}$
$=$ $\mathbb{E}\Phi\left(z_{\alpha}+T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}\right)\mathbb{I}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|\leq\epsilon\}$
$\geq$ $\mathbb{E}\Phi\left(z_{\alpha}+T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}\right)$
$\qquad-\mathbb{E}\Phi\left((z_{\alpha}+T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}\right)\mathbb{I}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|>\epsilon\}$
$\geq$ $\mathbb{E}\Phi\left(z_{\alpha}+T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}\right)-\mathbb{P}(|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|>\epsilon),$
where the first inequality is obtained due to $\Phi$ is non-negative and the last inequality holds because $T_{w}^{(2)}(\bm{X})\sigma_{q,L}>0$ under stochastic dominance assumption (Assumption 3) and $\Phi$ is bounded by $1$. Together with Lemma S2 and Assumption 4, we obtain
$\begin{split}&\mathbb{P}_{\bm{X}\sim p}\left(T_{w}(\bm{X})\leq z_{\alpha}\right)\\ \geq&\min\left\{1-\alpha,\alpha+\phi(z_{\alpha})\mathbb{E}\left\{T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}}\right\}\right\}\frac{\widetilde{\sigma}_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}\\ &-\mathbb{P}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|\geq\epsilon\}+O\left(\log L/\sqrt{L}\right)+o(1).\end{split}$ (30)
Let $L\to\infty$ and using the fact that $\mathbb{E}\left\{T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}}\right\}=T_{w}^{(2*)}(\bm{X})$, we obtain that TNR is asymptotically lower bounded by $\min\{1-\alpha,\alpha+\phi(z_{\alpha})T_{w}^{(2*)}(\bm{X})\}\frac{\widetilde{\sigma}_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}$. By taking $\epsilon\to 0$, then the conclusion of Theorem 1 follows.
Remark 1.
It is worth noting that since $T_{w}^{(2)}(\bm{X})\sigma_{q,L}>0$, a sharper lower bound can be obtained by applying Lemma S3 in the last step (inequality (30)), then follow a similar argument, we will obtain that TNR is asymptotically lower bounded by
$\sup_{0<\beta\leq\alpha}\min\{1-\beta,\alpha+\phi(z_{\beta})T_{w}^{(2*)}(\bm{X})\}.$ (31)
Noted that with any fixed $\alpha\in(0,1/2)$, the lower bound in (31) may tend to $1$ given that $T_{w}^{(2*)}(\bm{X})$ is sufficiently large, which is a sharper bound than in Theorem 1.
E.4 Proof of Theorem 2
Proof.
Denote $Z_{t}=\widehat{w}\left(\log q_{t}(X_{t}|X_{<t})\right)-\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}(\bullet|X_{<t})}\widehat{w}\left(\log q_{t}(X_{t}|X_{<t})\right)$. Then if $\bm{X}\sim q$, we have $\mathbb{E}\left\{Z_{t}|X_{<t}\right\}=0$ almost surely. Under Assumptions 3, 5 and the boundedness of $\widehat{w}$ (the B-spline basis are bounded), it is easy to verify that $Z_{t}$ satisfies all conditions of Lemma S1. Therefore, according to Lemma S1, we obtain for any $\alpha\in(0,1)$,
$\text{FNR}_{\widehat{w}}-\alpha$ $=$ $\mathbb{P}_{\bm{X}\sim q}(T_{\widehat{w}}(\bm{X})\leq z_{\alpha})-\Phi(z_{\alpha})$
$=$ $\mathbb{P}_{\bm{X}\sim q}\left(\frac{\sum_{t=1}^{L}Z_{t}}{\sum_{t=1}^{L}\mathbb{E}\{Z_{t}^{2}|X_{<t}\}}\leq z_{\alpha}\right)-\Phi(z_{\alpha})$
$\leq$ $O\left(\frac{\log L}{\sqrt{L}}\right)+O\big((\mathbb{E}\big|V_{L}-1\big|)^{1/3}\big).$
Taking expectation on both sides, we obtain
$\mathbb{E}(\text{FNR}_{\widehat{w}})\leq\alpha+O\left(\frac{\log L}{\sqrt{L}}\right)+O\big((\mathbb{E}\big|V_{L}-1\big|)^{1/3}\big).$
This completes the proof. ∎
E.5 Proof of Theorem 3
Proof.
Since $\mathbb{E}(\text{TNR}_{\widehat{w}})\geq\text{TNR}_{w^{*}}-\mathbb{E}(|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}|)$, it is enough to upper bound the second term in the last expression. Denote by $\widehat{T}_{n}(\bullet)$ and $\widehat{T}^{*}(\bullet)$ respectively the classifier (4) using witness functions $\widehat{w}(\bullet)=\phi(\bullet)^{\top}\widehat{\beta}$ and $w^{*}(\bullet)=\phi(\bullet)^{\top}\beta^{*}$ (see the definition of $\widehat{\beta}$ and $\beta^{*}$ in Lemma S4). Write $\widehat{\Delta}_{n}(\boldsymbol{x})=|\widehat{T}_{n}(\boldsymbol{x})-T^{*}(\boldsymbol{x})|$ and $\widehat{\Delta}_{n}=\sup_{\boldsymbol{x}}\Delta_{n}(\boldsymbol{x})$. For any $z_{\alpha}>0$ we have that
$\left|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}\right|=$ $\left|\mathbb{P}_{\boldsymbol{X}\sim p}\left(\widehat{T}_{n}\left(\boldsymbol{X}\right)\leq z_{\alpha}\right)-\mathbb{P}_{\boldsymbol{X}\sim p}\left(T^{*}\left(\boldsymbol{X}\right)\leq z_{\alpha}\right)\right|$
$=$ $\left|\int\boldsymbol{1}_{\left\{\widehat{T}_{n}\left(\boldsymbol{x}\right)\leq z_{\alpha}\right\}}-\boldsymbol{1}_{\left\{T^{*}\left(x\right)>z_{\alpha}\right\}}\mathrm{d}p\left(\boldsymbol{x}\right)\right|$
$\leq$ $\int\left|\boldsymbol{1}_{\left\{\widehat{T}_{n}\left(\boldsymbol{x}\right)\leq z_{\alpha}\right\}}-\boldsymbol{1}_{\left\{T^{*}\left(x\right)>z_{\alpha}\right\}}\right|\mathrm{d}p\left(\boldsymbol{x}\right)$
$=$ $\int\boldsymbol{1}_{\left\{\widehat{T}_{n}(\boldsymbol{x})\leq z_{\alpha},T^{*}(\boldsymbol{x})>z_{\alpha}\right\}}+\boldsymbol{1}_{\left\{\widehat{T}_{n}(\boldsymbol{x})>z_{\alpha},T^{*}(\boldsymbol{x})\leq z_{\alpha}\right\}}\mathrm{d}p\left(\boldsymbol{x}\right)$
$\leq$ $2\int\boldsymbol{1}_{\left\{|T^{*}(\boldsymbol{x})-z_{\alpha}|\leq\widehat{\Delta}_{n}\right\}}\mathrm{d}p\left(\boldsymbol{x}\right)$
$=$ $2\mathbb{P}_{\boldsymbol{X}\sim p}\left(\left|T^{*}(\boldsymbol{X})-z_{\alpha}\right|\leq\widehat{\Delta}_{n}\right).$ (32)
Due to Assumption 1 on the event
$\left\{\widehat{\Delta}_{n}\leq\delta_{0}\right\}$ (33)
we will have that $\left|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}\right|\leq\kappa_{3}\widehat{\Delta}_{n}$ for some absolute $\kappa_{3}$. We therefore focus on bounding the quantity $\widehat{\Delta}_{n}$. For each $w\in\Omega$ and each $j=1,\dots,L$, we introduce the quantities:
$Y_{j}^{(w)}=w\left(\log q_{t}\left(X_{t}\mid X_{<t}\right)\right),$
$\mu_{j}^{(w)}=\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}w\left(\log q_{t}\left(\widetilde{X}_{t}\mid X_{<t}\right)\right),$
$(\sigma_{j}^{(w)})^{2}=\textup{Var}_{\widetilde{X}_{t}\sim q_{t}}w\left(\log q_{t}\left(\widetilde{X}_{t}\mid X_{<t}\right)\right).$
With this notation in place we have that for any $\boldsymbol{x}$
$\widehat{\Delta}_{n}(\boldsymbol{x})$ $\leq\frac{1}{\sqrt{L}}\left|\sum_{j=1}^{L}y_{j}^{(\widehat{w})}-\mu_{j}^{(\widehat{w})}\right|\left|\sqrt{L^{-1}\sum_{j=1}^{L}\left(\sigma_{j}^{(\widehat{w})}\right)^{2}}-\sqrt{L^{-1}\sum_{j=1}^{L}\left(\sigma_{j}^{(w^{*})}\right)^{2}}\right|^{-1}$ (34a)
$\hskip 20.00003pt+\left\{\frac{1}{L}\sum_{t=1}^{L}\left(\sigma_{j}^{(w^{*})}\right)^{2}\right\}^{-\frac{1}{2}}\frac{1}{\sqrt{L}}\left|\sum_{j=1}^{L}\left(Y_{j}^{(\widehat{w})}-Y_{j}^{(w^{*})}\right)-\left(\mu_{j}^{(\widehat{w})}-\mu_{j}^{(w^{*})}\right)\right|,$ (34b)
where for clarity we have suppressed dependence on $\boldsymbol{x}$ above. For ease of notation put $Z_{t}=\log q_{t}(X_{t}\mid X_{<t})$ and $\widetilde{Z}_{t}=\log q_{t}(\widetilde{X}_{t}\mid X_{<t})$ where $\widetilde{X}_{t}\sim q_{t}$. Write also $\phi(\bullet)=(B_{1}(\bullet),\dots,B_{d}(\bullet))^{\top}$. Recalling that $w(\bullet)=\phi(\bullet)^{\top}\beta$ for arbitrary $j=1,\dots,L$ we have
[…]
$\leq\frac{\kappa_{4}}{2}\sum_{l_{1}=1}^{d}\sum_{l_{2}=1}^{d}\left|\widehat{\beta}_{l_{1}}\widehat{\beta}_{l_{2}}-\beta_{l_{1}}^{*}\beta_{l_{2}}^{*}\right|$
$=\frac{\kappa_{4}}{2}\sum_{l_{1}=1}^{d}\sum_{l_{2}=1}^{d}\left|\widehat{\beta}_{l_{1}}\left(\widehat{\beta}_{l_{2}}-\beta_{l_{2}}^{*}\right)-\beta_{l_{2}}^{*}\left(\beta_{l_{1}}^{*}-\widehat{\beta}_{l_{1}}\right)\right|$
$\leq\frac{\kappa_{4}}{2}\sum_{l_{1}=1}^{d}\sum_{l_{2}=1}^{d}\left|\widehat{\beta}_{l_{1}}\right|\left|\widehat{\beta}_{l_{2}}-\beta_{l_{2}}^{*}\right|+\frac{\kappa_{4}}{2}\sum_{l_{1}=1}^{d}\sum_{l_{2}=1}^{d}\left|\beta_{l_{2}}^{*}\right|\left|\widehat{\beta}_{l_{1}}-\beta_{l_{1}}^{*}\right|$
$=\kappa_{5}\sqrt{d}\left\|\widehat{\beta}-\beta^{*}\right\|_{1}\kappa_{5}$
$\leq d\left\|\widehat{\beta}-\beta^{*}\right\|_{2},$
for absolute $\kappa_{4},\kappa_{5}$. Consequently, using inequality (23) and Assumption 2, on the event (27) we obtain that with absolute $\kappa_{6}$:
[…]
for certain absolute $\kappa_{7},\kappa_{8}$. Similar arguments show that the normalized sum in (34b) has the same tail behavior as (36). Since the above augments do not depend on $\boldsymbol{x}$ we obtain that (36) likewise described the tail behavior of $\widehat{\Delta}_{n}$. Consequently, on the relevant events we obtain that
$\mathbb{E}\left|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}\right|$ $=\int_{0}^{\infty}\mathbb{P}\left(\left|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}\right|>z\right)\mathrm{d}z$
$\leq\int_{0}^{\infty}\mathbb{P}\left(\kappa_{3}\widehat{\Delta}_{n}>z\right)\mathrm{d}z$
$\leq\kappa_{9}\int_{0}^{\infty}\exp\left(-\kappa_{10}\min\left\{z^{2}\frac{n}{d^{5\gamma+6}},z\sqrt{\frac{n}{d^{5\gamma+6}}}\right\}\right)\mathrm{d}z$
$\leq\kappa_{11}\sqrt{\frac{d^{5\gamma+6}}{n}}$ (37)
for certain absolute $\kappa_{9},\kappa_{10},\kappa_{11}$. When the events (33) and (23) do not hold from (32) we have the conservative bound $\mathbb{E}\left|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}\right|\leq 2$. However, the probability of these events not holding is smaller than (37) up to constants. Therefore, the stated result follows by the law of total expectation. ∎
Appendix F Experiment details
Pre-trained language models. We assess the performance of our method using text generated from various pre-trained language models outlined in Table S6. Following the setting in Bao et al. (2024), for the models with over 6B parameters, we employ half-precision (torch.float16), otherwise, we use full-precision (torch.float32).
Name Model Scale (Billion)
GPT-2 (Radford et al., 2019) openai-community/gpt2-xl 1.5B
GPT-Neo (Black et al., 2021) EleutherAI/gpt-neo-2.7B 2.7B
OPT-2.7 (Zhang et al., 2022) facebook/opt-2.7b 2.7B
GPT-J (Wang & Komatsuzaki, 2021) EleutherAI/gpt-j-6B 6B
Qwen2.5 (Yang et al., 2024a) Qwen/Qwen2.5-7B 7B
Mistral (Jiang et al., 2023) mistralai/Mistral-7B-v0.3 7B
Llama (AIMeta, 2024) meta-llama/Meta-Llama-3-8B 8B
GPT-NeoX (Black et al., 2022) EleutherAI/gpt-neox-20b 20B
Table S6: Description of the source models that is used to produce machine-generated text. : we present the address of models in https://huggingface.co/.
Setup of the closed-source LLMs. For the gpt-4o, the version is set as gpt-4o-2024-08-06. The generation process by sending the following messages to the service. Message for XSum and Writing is the same as that described in Section C.2 in Bao et al. (2024). We describe that for Yelp and Essay are:
{'role': 'system', 'content': 'You are a Review writer on Yelp.'},
{'role': 'user', 'content': 'Please write an article with about 150 words starting exactly with: <prefix>'},
and
{'role': 'system', 'content': 'You are a student of high school and university level. And now, you are an Essay writer.'},
{'role': 'user', 'content': 'Please write an essay with about 200 words starting exactly with: <prefix>'},
respectively.
For Claude-3.5-Haiku, the system instruction was set analogously (e.g., Yelp review or essay writer), while the user role contained the corresponding content prompt.
For Gemini, the instruction was fed into the system_instruction parameter, with a value identical to the concatenation of the system content and the user content used for GPT-4o.
For all closed-source models, the temperature parameter is set to 0.8 to encourage the generated text to be creatively diverse and less predictable.
Implementations of baselines. For the baselines considered in our experiments, we use the existing implementation provided in https://github.com/baoguangsheng/fast-detect-gpt, which is distributed in the MIT License. We run DetectGPT and NPR with default 100 perturbations with the T5 model (Raffel et al., 2020) and run DNA-GPT with a truncate-ratio of 0.5 and 10 prefix completions per passage.
Evaluation Metric. We measure the detection accuracy by AUC (short for “area under the curve”). AUC ranges from 0.0 to 1.0, an AUC of 1.0 indicates a perfect classifier and vice versa. The relative improvement of AdaDetectGPT over FastDetectGPT is calculated by $\frac{\textup{AdaDetectGPT}-\textup{FastDetectGPT}}{1.0-\textup{FastDetectGPT}}$, which represents how much improvement has been made relative to the maximum possible improvement for FastDetectGPT.
Hardware details. Most of experiments are run on a Tesla A100 GPU (40GB) with 10 vCPU Intel Xeon Processor and 72GB RAM. For the experiments where the source model is GPT-NeoX, we run on a H20-NVLink (96GB) GPU with 20 vCPU Intel(R) Xeon(R) Platinum and 200GB RAM.
Appendix G Additional numerical results
G.1 Results on additional open-source models
Model Method XSum Writing Essay SQuAD Yelp
Qwen2.5 Likelihood 0.6175 0.7041 0.6755 0.5183 0.6793
Entropy 0.5403 0.5043 0.5073 0.5232 0.5236
LogRank 0.6325 0.7150 0.6958 0.5166 0.6943
Binoculars 0.6297 0.7578 0.8018 0.6164 0.7199
TextFluoroscopy 0.5778 0.5110 0.5638 0.5383 0.5060
RADAR 0.6469 0.6190 0.6061 0.6262 0.6276
ImBD 0.6653 0.6584 0.7874 0.5168 0.7392
BiScope 0.6320 0.6610 0.6625 0.6250 0.7050
Fast-DetectGPT 0.7523 0.8513 0.8347 0.5016 0.8465
AdaDetectGPT 0.7963 0.8965 0.8799 0.6044 0.8915
Relative 17.7682 30.3912 27.3167 0.6431 29.3165
Mistral Likelihood 0.7409 0.8643 0.8667 0.7068 0.7598
Entropy 0.5290 0.5420 0.6052 0.5070 0.5103
LogRank 0.7270 0.8446 0.8467 0.7041 0.7499
Binoculars 0.7218 0.8440 0.8314 0.7258 0.7502
TextFluoroscopy 0.6210 0.5555 0.5127 0.5772 0.5109
RADAR 0.6518 0.6537 0.6292 0.6055 0.6018
ImBD 0.7683 0.8391 0.8631 0.8073 0.7440
BiScope 0.7320 0.8740 0.9000 0.7283 0.7840
Fast-DetectGPT 0.8922 0.8151 0.9052 0.8812 0.8902
AdaDetectGPT 0.8944 0.8275 0.9069 0.8851 0.9026
Relative 2.0423 6.6718 1.8332 3.3051 11.2763
LLaMA3 Likelihood 0.8236 0.8929 0.9115 0.7071 0.8915
Entropy 0.5545 0.5732 0.5626 0.5047 0.5010
LogRank 0.8634 0.9122 0.9351 0.7422 0.9146
Binoculars 0.9546 0.9845 0.9949 0.9469 0.9854
TextFluoroscopy 0.5479 0.5274 0.5478 0.5535 0.5362
RADAR 0.7154 0.7285 0.7835 0.6619 0.7875
ImBD 0.8643 0.8837 0.8928 0.7596 0.8677
BiScope 0.9450 0.9830 0.9900 0.8783 0.9860
Fast-DetectGPT 0.9734 0.9879 0.9901 0.9488 0.9882
AdaDetectGPT 0.9782 0.9893 0.9924 0.9553 0.9900
Relative 18.0119 11.6202 22.7215 12.6288 15.7610
Table S7: Performance on three open-source LLMs (Qwen2.5, Mistral, LLaMA3) across five datasets.
G.2 Computational analysis
From Table S8, the runtime of AdaDetectGPT is around 44 seconds and changes marginally with respect to $d$. This is because we can use a closed-form expression to learn a witness function. This time is nearly negligible compared with the time required to compute logits, which involves feeding tokens from multiple passages into LLMs. Furthermore, the training time when $n$ increases is shown in Table S9. It shows AdaDetectGPT saves more runtime and memory than ImBD.
$d$ 4 8 12 16 20
Runtime 44.48 44.62 44.73 44.92 45.00
Table S8: Runtime scale with $d$.
$n$ 100 150 200 250 300 350
ImBD 45.65(9.56) 68.20(9.57) 93.00(9.57) 116.08(9.56) 136.53(9.56) 158.15(9.58)
AdaDetectGPT 9.28(0.36) 23.45(0.37) 40.19(0.37) 44.25(0.37) 59.57(0.60) 69.56(0.37)
Table S9: Runtime (memory in parenthesis) scale with $n$ comparing with ImBD. The runtime is measured in seconds and the memory is measured in GB.
G.3 Analysis factors in training $w$
Since AdaDetectGPT requires training a witness function, we systematically examine three key factors influencing its performance: (1) the size of the training set; (ii) tuning parameters for generating B-spline basis and (iii) distribution shift between training and test data.
Robust performance on various training data sizes. We evaluate AdaDetectGPT across varying dataset sizes by setting $n_{1}=n_{2}\in\{100,200,300,400,500,600\}$ for both human- and machine-generated texts. Figure S5 demonstrates that AdaDetectGPT has a clear performance advantage over FastDetectGPT when sample size is large. This is expected because a larger sample size leads to a more accurate estimation of $w$. Notably, even with limited data $n_{1}=n_{2}=100$, AdaDetectGPT maintains superior accuracy compared to baseline methods, though the performance gap decreases. These results highlight our method’s effectiveness on learning the witness function.
Insensitivity on tuning parameters. The two critical tuning parameters of B-splines are: (i) the number of basis functions (n_base) and (ii) the maximum polynomial order. Our experiments fix one parameter while varying the other (with n_base=16 or order=2 as defaults). As shown in Figure S6 in Appendix, AdaDetectGPT achieves the highest AUC scores so long as n_base $\geq 4$. Besides, enlarging n_base improves the AUC of AdaDetectGPT although the improvement becomes marginal when n_base $\geq 16$. Figure S6 also shows that increasing the polynomial order from linear to quadratic visibly improve the performance; while increasing order from quadratic to cubic/quartic has a limited gain. Having said that, even when the B-splines function is piecewise linear, our method still outperform all baselines.
Robust against distribution shift. In this part, we create training sets with different distributions than the test data by varying the number of human prompt tokens in machine-generated text. In contrary, for the test data, the number of human prompt tokens are fixed. As shown in Figure S7, AdaDetectGPT demonstrates high robustness - the distributional discrepancy between training and test data has negligible effect on classification accuracy. Notably, our method always achieves the highest AUC across all experimental setup.
Figure S5: Classification accuracy versus the sample size for training $w$. We omit DetectGPT, NPR, and DNA in this experiments as they are time-consuming.
Figure S6: The classification accuracy of AdaDetectGPT and baseline methods. AdaDetectGPT(n_base) present the AUC when the number of basis in B-spline increases as 4, 8, 16, 32, 64 (bottom $x$-axis); while AdaDetectGPT(order) shows the AUC when the maximum order of basis in B-spline increases from 1 to 4 (top $x$-axis). The AUC of baseline methods are presented by dash lines.
Figure S7: The classification accuracy of AdaDetectGPT when the number of human prompts changes. The AUC of baseline methods are presented by dash lines.
G.4 Robustness on adversarial Attacks
We have conducted additional experiments to evaluate the robustness of AdaDetectGPT against 2 adversarial attacks: (i) paraphrasing, where an LLM is instructed to rephrase human-written text, and (ii) decoherence, where the coherence LLM-generated text is intentionally reduced to avoid detection. These experiments were carried out across 3 datasets and 3 types of sampling and scoring models setup, resulting in a total of 18 settings. Both adversarial attacks were implemented following Bao et al. (2024).
paraphrasing decoherence
DetectGPT Xsum Writing PubMed Avg. Xsum Writing PubMed Avg.
Fast (GPT-J/GPT-2) 0.9178 0.9137 0.7944 0.8753 0.7884 0.9595 0.7870 0.8449
Ada (GPT-J/GPT-2) 0.9225 0.9121 0.8029 0.8792 0.8765 0.9597 0.8284 0.8882
Fast (GPT-J/Neo-2.7) 0.9602 0.9185 0.7310 0.8699 0.8579 0.9701 0.7609 0.8630
Ada (GPT-J/Neo-2.7) 0.9623 0.9181 0.7587 0.8797 0.9230 0.9704 0.8124 0.9019
Fast (GPT-J/GPT-J) 0.9537 0.9458 0.7041 0.8679 0.8836 0.9869 0.7550 0.8752
Ada (GPT-J/GPT-J) 0.9587 0.9449 0.7308 0.8781 0.9336 0.9864 0.8008 0.9070
Table S10: Detection of LLM texts under two adversarial attacks.
Results are reported in the Table S10. It can be seen that, AdaDetectGPT consistently achieves higher AUCs than Fast-DetectGPT in most of the 18 scenarios. And the improvement reaches up to 10% for paraphrasing and up to 85% for decoherence. These results suggest that AdaDetectGPT attains more robust performance against adversarial attacks.
G.5 Accuracy on open-source models under black-box setting
Dataset Method Source Model
GPT-2 OPT-2.7 GPT-Neo GPT-NeoX Avg.
SQuAD FastDetectGPT 0.6181 0.6495 0.6230 0.6910 0.6813
AdaDetectGPT 0.6920 0.7195 0.7382 0.7338 0.7460
Relative 19.3570 19.9651 30.5609 13.8495 20.2957
FastDetectGPT 0.8145 0.8166 0.9220 0.7519 0.8188
AdaDetectGPT 0.8249 0.8308 0.9273 0.7609 0.8300
Relative 5.6301 7.7245 6.7968 3.6121 6.2106
Writing FastDetectGPT 0.7662 0.7918 0.7685 0.8022 0.8028
AdaDetectGPT 0.8306 0.8529 0.8555 0.8587 0.8636
Relative 27.5699 29.3365 37.6112 28.5350 30.8124
FastDetectGPT 0.8565 0.8497 0.9215 0.8182 0.8582
AdaDetectGPT 0.8780 0.8737 0.9386 0.8567 0.8849
Relative 14.9666 15.9741 21.7742 21.1856 18.8023
XSum FastDetectGPT 0.5919 0.6445 0.5718 0.6389 0.6468
AdaDetectGPT 0.6795 0.7238 0.6879 0.7045 0.7261
Relative 21.4569 22.2991 27.1129 18.1580 22.4439
FastDetectGPT 0.8145 0.8166 0.9220 0.7519 0.8188
AdaDetectGPT 0.8249 0.8308 0.9273 0.7609 0.8300
Relative 9.8060 10.5637 10.2543 8.1057 11.1574
Table S11: Zero-shot detection accuracy on five source models under the black-box setting. : use two surrogate models for sampling and scoring, where the sampling model is GPT-J while the scoring model is GPT-Neo.
Appendix H Broader impact and limitation
AdaDetectGPT is a statistically efficient detector for machine-generated text. By accurately detecting generated passages, it can help safeguard AI systems against fake news, disinformation, and academic plagiarism.
In this paper, we prove rigorous guarantees in the white-box setting and demonstrate empirical performance in the black-box setting. The promising results under black-box setting motivate future work to establish matching theoretical guarantees. Moreover, even in the white-box setting, practical LLM’s text generation often employs sampling parameters (e.g. temperature and top_k) that makes the sampling distribution differs from the conditional distribution derived from the source model. This mismatch causes MCLT may not hold in practice. Fortunately, we observe that our test statistic still shifts toward a positive mean (see Figure S8), implying FNR remains control.
Figure S8: Histogram of statistics in three dataset. Each panel visualizes the histogram of statistics. The yellow histogram corresponds to the case when the sampled texts exactly follow the conditional probability of the source model, while purple histogram corresponds to text drawn with deviations from that distribution.